### v1

# Deciphering interventional dynamical causality from non-intervention systems

[…]
-3pt plus 3pt minus 2pt
Deciphering interventional dynamical causality from non-intervention systems
Jifan Shi1,2,3, Yang Li4, Juan Zhao5, Siyang Leng1,6, Kazuyuki Aihara4∗, Luonan Chen4,7,8∗, Wei Lin1,2,3∗
1Research Institute of Intelligent Complex Systems, Fudan University, Shanghai 200433, China
2Shanghai Artificial Intelligence Laboratory, Shanghai 200232, China
3State Key Laboratory of Medical Neurobiology and MOE Frontiers Center for Brain Science, Fudan University, Shanghai 200032, China
4International Research Center for Neurointelligence, The University of Tokyo Institutes for Advanced Study, The University of Tokyo,Tokyo 113-0033, Japan
5School of Pharmacy, Shanghai University of Traditional Chinese Medicine, Shanghai 201203, China
6Academy for Engineering and Technology, Fudan University, Shanghai 200433, China
7Key Laboratory of Systems Biology, Shanghai Institute of Biochemistry and Cell Biology, Center for Excellence in Molecular Cell Science, Chinese Academy of Sciences, Shanghai 200031, China
8Key Laboratory of Systems Health Science of Zhejiang Province, School of Life Science, Hangzhou Institute for Advanced Study, University of Chinese Academy of Sciences, Chinese Academy of Sciences, Hangzhou 310024, China
To whom correspondence should be addressed; E-mail: kaihara@g.ecc.u-tokyo.ac.jp or lnchen@sibs.ac.cn or wlin@fudan.edu.cn.
Detecting and quantifying causality is a focal topic in the fields of science, engineering, and interdisciplinary studies. However, causal studies on non-intervention systems attract much attention but remain extremely challenging. To address this challenge, we propose a framework named Interventional Dynamical Causality (IntDC) for such non-intervention systems, along with its computational criterion, Interventional Embedding Entropy (IEE), to quantify causality. The IEE criterion theoretically and numerically enables the deciphering of IntDC solely from observational (non-interventional) time-series data, without requiring any knowledge of dynamical models or real interventions in the considered system. Demonstrations of performance showed the accuracy and robustness of IEE on benchmark simulated systems as well as real-world systems, including the neural connectomes of C. elegans, COVID-19 transmission networks in Japan, and regulatory networks surrounding key circadian genes.
Introduction
Directional or indirectional interactions among different components give rise to a variety of complex phenomena in nature and social society. The causality or the causal effect is one of the most attractive directional relations originating from the non-reversibility of time. Modeling, detecting, and quantifying causality from observational data are crucial for describing, interpreting, predicting, and even controlling complex systems.
Statisticians believe that the causality is contained and obtainable in universal random variables regardless of the time label. The Neyman-Rubin’s potential outcome framework  (?, ?, ?), the Wright’s structural equation model  (?, ?) and the Pearl’s causal diagram model  (?, ?) are the most famous statistical approaches, which have been proved to be of mathematical equivalence  (?). Especially, the instrumental variable method, that was awarded the 2021 Nobel Prize in economic sciences, is widely applied in the study of causality among economics, environments, and other disciplines  (?). Statistical methods seek to discern binary causal relations among random variables on a directed acyclic graph, without relying on temporal data. However, in general complex dynamical systems, the causation must precede the effect, feedback is common, and the causal strength needs to be quantitatively measured. Time should play definitely a key role.
Dynamics-oriented researchers have proposed fruitful algorithms for measuring the observational causality from time series, commonly referred to as the dynamical causality  (?). The celebrated Granger causality (GC)  (?, ?) employs a linear model and uses improvement of predictability over time to illustrate the causality. Transfer entropy (TE)  (?) generalizes GC to the nonlinear case by quantifying the prediction uncertainty through Shannon entropy. Neither GC nor TE addresses the “non-separability” problem, which means that removing the causal variable from the system inevitably influences the dynamics of downstream variables  (?, ?, ?, ?). To measure causality in non-separable systems, numerous approaches have emerged in the last decade within the framework of delay embedding. These include convergence of cross mapping (CCM)  (?, ?), partial cross mapping (PCM)  (?), continuity scaling  (?), inverse continuity  (?), topological expansion  (?), joint distance distribution  (?), embedding entropy (EE)  (?) and other indices  (?, ?).
Nonetheless, most indices designed for quantifying the dynamical causality primarily estimate directional causal relations at the observational level, i.e. observational dynamical causality (ObsDC). Conventionally, to detect and quantify causality at the interventional level, which is noted as the interventional dynamical causality (IntDC) in this study, intelligent modulation and manipulation of the dynamical system should be taken into account. By allowing external intervention to the system and recording data under different perturbations, frameworks such as perturbation cascade inference (PCI)  (?) and dynamical causal effect (DCE)  (?, ?) provide relevant computational schemes. However, due to ethical or practical limitations, many real systems should be analyzed without any external intervention. An essential problem is how to measure IntDC solely from the observational data but at the interventional level.
In this study, we propose the IntDC framework and introduce a criterion named Interventional Embedding Entropy (IEE), which aims to identify and quantify IntDC between variables solely from the observational time series. Actually, IEE is rigorously derived by the theory of the delay embedding. Numerical IEE does not require specific prior knowledge of dynamics, and additional perturbation to the system is also unnecessary. Compared to ObsDC indices, IEE designed for IntDC has the capability to rank the importance of causal effects and construct directional causal networks more effectively. We demonstrate numerical experiments on both simulated examples and real datasets, including estimating neural connectomes of C. elegans, evaluating COVID-19 transmission in Japan, and constructing regulatory networks surrounding key circadian genes.
Methods
Observational dynamical causality in delay-embedding space
We consider a discrete two-variable dynamical system as
$\begin{cases}x_{t+1}=g(x_{t},x_{t-1},\dots,x_{t-p},\varepsilon_{x,t}),\\ y_{t+1}=f(x_{t},x_{t-1},\dots,x_{t-p},y_{t},y_{t-1},\dots,y_{t-p},\varepsilon_{y,t}),\end{cases}$ (1)
where $x,y$ are two variables, $p$ denotes the time step during which causality is considered, and $\varepsilon_{\cdot,t}$ stand for small noise terms. According to Shi et al.  (?), there exists ObsDC from $x$ to $y$ since the evolving equation of $y_{t+1}$ depends on the historical behavior of $x$, while there is no ObsDC from $y$ to $x$ as the dynamics of $x_{t+1}$ is independent of $y$. GC and TE measure the predictability of $f$ in the original time-series space to quantify the ObsDC from $x$ to $y$ through linear regression and entropy uncertainty, respectively (Fig. 1(B)). The causal strength of GC/TE is based on the assumption of separability, which means that removing $x$ from the system does not influence the observed data of $y$. However, in general coupled systems, the separability is not satisfied, as removing one variable changes the values of affected ones  (?, ?, ?).
To address the challenge of detecting ObsDC in universal non-separable systems, various approaches from the delay-embedding space have been proposed. In the delay-embedding framework, the bivariate system Eq. (1) is assumed to evolve into an attractive manifold with an inner dimension $d$. Let the time-delay vectors of $x$ and $y$ be
$\bm{X}_{t}=(x_{t},x_{t-1},\dots,x_{t-L})^{T}\in\mathcal{M}_{X}\subseteq\mathbb{R}^{L+1},$ (2)
and
$\bm{Y}_{t+1}=(y_{t+1},y_{t},y_{t-1},\dots,y_{t-L+1})^{T}\in\mathcal{M}_{Y}\subseteq\mathbb{R}^{L+1},$ (3)
where $L$ is the time-delay length, $\mathcal{M}_{X}$ and $\mathcal{M}_{Y}$ represent the manifolds formed by $\bm{X}_{t}$ and $\bm{Y}_{t}$, respectively. According to the seminal stochastic version of Takens’ embedding theorem  (?, ?, ?, ?), for delay dimension $L\geqslant 2d$, we can obtain
$\bm{X}_{t}=\bm{F}(\bm{Y}_{t+1}),$ (4)
where $\bm{F}$ is a smooth projection operator. The detailed derivation can be referred to the Supplementary Text. Equation (4) shows that the causal variable $\bm{X}_{t}\in\mathcal{M}_{X}$ can be reconstructed by the effect variable $\bm{Y}_{t+1}\in\mathcal{M}_{Y}$. However, as $\bm{F}$ is not reversible in generic sense, we can not determine $\bm{Y}_{t+1}$ only by the information from $\bm{X}_{t}$. The causal dependence in $f$ between $x$ and $y$ is transformed into the reconstructability of $\bm{F}$ (Fig. 1(D)). Instead of fitting a model by removing the causal variable as in GC and TE, detecting the existence and quantifying the continuity characteristics of $\bm{F}$ in the delay-embedding space can be adequate for causal identification, especially for universal non-separable systems. Related algorithms include CCM, PCM, and EE.
Figures 1(A), (B) and (D) summarize dynamical causality at the observational level, where the ObsDC from variable $x$ to $y$ is estimated only from the historical observed time series. But fitting Eq. (1) or Eq. (4) from data is criticized as mere association or prediction  (?, ?).
Figure 1: Illustration for the observational dynamical causality (ObsDC) and interventional dynamical causality (IntDC). (A) An illustrative example of a complex system where $(x_{t},y_{t})$ are observed time series. The $z_{t}$ stands for the unobserved variable. (B) ObsDC from $x$ to $y$ indicates that the dynamics of $y$ depends on $x$. Algorithms such as Granger causality (GC) and transfer entropy (TE) infer the ObsDC in the original time-series space. (C) IntDC from $x$ to $y$ indicates that the deviation of $y$, i.e. $\delta y$, depends on an intervention on $x$, i.e. $\delta x$. Usually additional data from a perturbed system $(\tilde{x}_{t},\tilde{y}_{t})$ is necessary for detecting IntDC. (D) ObsDC can be detected in the embedding space by cross mapping $\mathbb{F}$. The causal vector $X_{t}$ can be reconstructed by the effect vector $\bm{Y}_{t+1}$. (E) IntDC can be modeled in the embedding space. The interventional causal vector $\delta\bm{X}_{t}$ should be reconstructable by the effect vector $\delta\bm{Y}_{t+1}$ around $\bm{Y}_{t+1}$. The $\delta\bm{Y}_{t+1}$ is approximated by the distance from $\bm{Y}_{t+1}$ to its neighbors $\bm{Y}_{t_{k}+1}$, while $\delta\bm{X}_{t}$ is obtained by $X_{t_{k}}$ with the same time label $t_{k}$. (F) Interventional embedding entropy (IEE) measures the average information retention contained in $\delta\bm{Y}_{t+1}$ from $\delta\bm{X}_{t}$, quantifying the IntDC from variable $x$ to $y$. Notably, IEE does not require additional perturbations to the system and can be inferred from observational data solely.
Interventional dynamical causality in delay-embedding space
The detection of hidden essential-level causality necessitates intervention or manipulation to the system  (?, ?, ?, ?). We define that there is an IntDC from $x$ to $y$ during the time $[t-p,t]$ in the original time-series space, if $\delta y_{t}=\tilde{y}_{t}-y_{t}$ depends on $\delta x_{t}=\tilde{x}_{t}-x_{t}$, where the dynamics of $x_{t}$ is perturbed to $\tilde{x}_{t}$, and $\tilde{y}_{t}$ is the dynamics of affected $y$ after the perturbation (Fig. 1(C)). The intervention $\delta x_{t}$ can be either a pulse stimulation or persistent disturbances to the system.
To measure the IntDC, data before and after the intervention, i.e. $(x_{t},y_{t})$ and $(\tilde{x}_{t},\tilde{y}_{t})$, are typically required, as in PCI and DCE. These conventional algorithms rely on reproducible dynamics or data availability under different settings.
To address challenges of causal inference in widespread non-intervention, non-linear and non-separable systems, we turn to the study of intervened dynamics Eq. (4) in the delay-embedding space. Denote $\delta\bm{X}_{t}=\widetilde{\bm{X}}_{t}-\bm{X}_{t}$ and $\delta\bm{Y}_{t+1}=\widetilde{\bm{Y}}_{t+1}-\bm{Y}_{t+1}$, where $\widetilde{\bm{X}}_{t}$ and $\widetilde{\bm{Y}}_{t+1}$ are the time-delay vectors of $x$ and $y$ after the intervention, respectively. Under infinitesimal intervention, Eq. (4) leads to
$\delta\bm{X}_{t}=\nabla_{\bm{Y}}\bm{F}(\bm{Y}_{t+1})\cdot\delta\bm{Y}_{t+1}.$ (5)
Equation (5) indicates that $\delta\bm{Y}_{t+1}$ contains complete information of $\delta\bm{X}_{t}$ in the neighborhood of $\bm{Y}_{t+1}$ or reconstructs $\delta\bm{X}_{t}$, if there exists IntDC from the causal variable $x$ to the effect variable $y$ (Fig. 1(E)).
Figures 1(A), (C), and (E) give an overview of the IntDC framework.
Interventional embedding entropy
We introduce the IEE criterion, specifically designed for quantifying IntDC solely from the observational data in the delay-embedding space. The IEE criterion is proposed to measure the average information retention or reconstructability from the effect $\delta\bm{Y}_{t+1}$ to the causation $\delta\bm{X}_{t}$, formulated as
$\text{IEE}[x\rightarrow y]=\text{CMI}(\delta\bm{X}_{t},\delta\bm{Y}_{t+1}|\bm{Y}_{t+1}),$ (6)
where CMI is the conditional mutual information  (?), i.e.
$\text{CMI}(\bm{x},\bm{y}|\bm{z})=\iiint p(\bm{x},\bm{y},\bm{z})\log\frac{p(\bm{x},\bm{y}|\bm{z})}{p(\bm{x}|\bm{z})p(\bm{y}|\bm{z})}\mathop{}\!\mathrm{d}\bm{x}\mathop{}\!\mathrm{d}\bm{y}\mathop{}\!\mathrm{d}\bm{z}.$ (7)
The $\delta\bm{Y}_{t+1}$ is estimated by the distance from $\bm{Y}_{t+1}$ to its neighbored points $\bm{Y}_{t_{k}+1}$ in the embedding space, and $\delta\bm{X}_{t}$ is calculated by $\bm{X}_{t_{k}}$ with the same time label $t_{k}$ (Figs. 1(E) and (F)). Thus, Eq. (6) can be numerically approximated only from the observational data, especially for non-intervention systems. Table 1 shows the numerical algorithm for IEE, whose detailed derivation can be found in the Supplementary Text.
Results
IEE can accurately and robustly quantify causality
To assess the numerical performance of measuring IntDC, we first applied the IEE algorithm on the following two-node Logistic dynamics
$\begin{cases}x_{t+1}=3.7\big{[}(1-\beta_{yx})x_{t}\left(1-x_{t}\right)+\beta_{yx}y_{t}\left(1-y_{t}\right)\big{]}+\varepsilon_{x,t},\\ y_{t+1}=3.7y_{t}\big{[}1-(1-\beta_{xy})y_{t}-\beta_{xy}x_{t}\big{]}+\varepsilon_{y,t},\end{cases}$ (8)
where $\varepsilon_{\cdot,t}$ represents independent Gaussian noises, and parameters $\beta_{xy}$ and $\beta_{yx}$ modulate the IntDC strength from $x$ to $y$ and from $y$ to $x$, respectively (see further details in SM).
IEE can accurately capture the numerical behavior of dynamical causality in the two-node Logistic system, especially IntDC rather than ObsDC. In Fig. 2(A), we set $\beta_{xy}\equiv 0$ and let $\beta_{yx}$ increase from 00 to $0.3$. We observed a monotonic increase in $\text{IEE}[y\rightarrow x]$ (the red line with squares), while $\text{IEE}[x\rightarrow y]$ remained around zero (the blue line with dots). In Fig. 2(B), we set $\beta_{xy}\equiv 0.1$ and let $\beta_{yx}$ increase from 00 to $0.3$. $\text{IEE}[y\rightarrow x]$ exhibited a monotonic increase (the red line with squares), while $\text{IEE}[x\rightarrow y]$ stayed above zero (the blue line with dots). The gray dashed lines in Figs. 2(A-B) represent $0.01$ for reference. A detailed comparison with GC, TE and CCM can be found in figs. S1 and S2, which indicate that GC exhibited non-monotonic behavior, $\text{TE}[y\rightarrow x]$ slightly decreased with a false-negative issue, and $\text{CCM}[x\rightarrow y]$ had a false-positive problem. The mean values and standard deviations of the indices under 100 simulations are provided in tables S1-S4.
Within suitable parameter ranges, IEE demonstrates stability and robustness. In Figs. 2(D-G), we validated the robustness of the IEE algorithm under various conditions, including different delayed lags $L$, the numbers of nearest neighbors $K$, lengths of time series $N$, and noise amplitudes $\sigma$. The causal strength should exhibit a descending trend with decreasing $\beta_{yx}$ from $0.15$ to $0.05$ through $0.125,0.1,0.075$. The experiments support that IEE maintains the correct order under different parameter values with relatively stable variations.
IEE can eliminate the effects of confounding variables
IEE remains applicable and accurate even in presence of confounding variables. We simulated a three-node Logistic dynamics $(x_{t},y_{t},z_{t})$, where $z$ acted as a confounder (see details in SM). There were consistently non-zero causal effects from $z$ to $x$ and from $z$ to $y$, while $\beta_{xy}$ adjusted the causal strength from $x$ to $y$ (Fig. 2(C)). When $\beta_{xy}=0$, $\text{IEE}[x\rightarrow y]$ was almost zero and significantly different from $\text{IEE}[z\rightarrow x]$ and $\text{IEE}[z\rightarrow y]$ (Fig. 2(C)); but conventional ObsDC indices such as GC/TE/CCM produced false-positive causality from $x$ to $y$ due to the presence of confounder $z$ (fig. S3). That is because an intervention on $x$, i.e. $\delta x$, will not induce $\delta y$ through $z$ under the IntDC framework. IEE, specially designed for measuring IntDC, remained unaffected by the confounder $z$ and yielded accurate results.
Figure 2: Performance of IEE on the Logistic systems. (A) and (B) used the two-node Logistic system. The $\beta_{xy}$ and $\beta_{yx}$ control the coupling coefficients from $x$ to $y$ and from $y$ to $x$, respectively. (A) shows the mean values of IEE when $\beta_{xy}=0$ and $\beta_{yx}$ ranges from 00 to $0.3$ over 100 simulations. $\text{IEE}[y\rightarrow x]$ (the red line with squares) increases monotonically, while $\text{IEE}[x\rightarrow y]$ (the blue line with dots) stays around zero. (B) is under $\beta_{xy}=0.1$, in which $\text{IEE}[x\rightarrow y]$ (the blue line with dots) is significantly positive. The gray dashed lines in (A) and (B) represent the constant $0.01$ as reference. (C) shows the performance of IEE in the three-node Logistic system, where $z$ acts as a confounding variable. When $\beta_{xy}=0$, $\text{IEE}[x\rightarrow y]$ is almost zero correctly. IEE designed for measuring IntDC is not affected by confounders, which usually leads to false positives. The results for $\beta_{xy}=0.5$ is shown for a comparison. The gray dashed line is $0.05$ for reference. (D-G) demonstrate the robustness of IEE under different delayed lags $L$, the numbers of nearest neighbors $K$, data lengths $N$, and noise amplitudes $\sigma$ in the two-node logistic system, respectively. The parameter $\beta_{xy}\equiv 0$, and $\beta_{yx}$ takes values of $0.15$ (red), $0.125$ (blue), $0.1$ (black), $0.075$ (green), and $0.05$ (magenta). Lines denote mean values under 50 simulations, and shaded areas represent the standard deviation. IEE maintains correct orders of IntDC under different situations with relatively stable variations.
IEE can reconstruct causal networks with quantified causal strength
IEE has the capability to reconstruct the causal network structure and rank the causal influence between nodes in complex networks at the interventional level. We used a 10-node coupled Henon maps as an example, where each node serves as the dynamical cause to its subsequent node (Fig. 3(A)). Further details regarding the dynamics can be found in SM.
To verify the effectiveness of IEE in reconstructing causal networks, we calculated the Area Under Curve (AUC) values based on multiple simulated time series. The AUC value of IEE was $0.871\pm 0.008$, significantly higher than conventional ObsDC indices, such as GC with $0.707\pm 0.069$, TE with $0.816\pm 0.018$, and CCM with $0.837\pm 0.012$. The boxplots for AUC values are displayed in Fig. 3(B). AUCs for IEE exhibited a higher mean and a smaller variation compared to the others. Additionally, fig. S4 presents receiver operating characteristic (ROC) curves.
To show the capability for ranking the importance of causal influences, we conducted two tests. We first calculated the IEE from Node 1 to the other nine nodes. Consistently with the actual scenario, IEE displayed sequential decrease in IntDC originating from Node 1 across Nodes 2 to 10 (Fig. 3(C)). Then, we measured the IEE received by Node 7 from the other nine nodes. As expected, Nodes 1-6 exhibited ascending causal strengths on Node 7, whereas Nodes 8-10 had little influence on Node 7 (Fig. 3(D)). IEE accurately discerned the IntDCs, with Node 6 showing the strongest value. A comparison between IEE for the IntDC and the results from GC, TE, and CCM for the ObsDC can be found in figs. S5 and S6. GC and TE failed to rank the influence from Node 1 accurately (fig. S5), while CCM suffered from the false-positive causal detection from Node 8 to Node 7 (fig. S6).
Figure 3: Performance of IEE on the 10-node coupled Henon-map dynamics. (A) illustrates the network structure of the dynamics. There are directed causality from each node to its subsequent node. (B) presents a boxplot of the Area Under Curve (AUC) values for the network reconstruction. IEE outperforms GC, TE, and CCM significantly in accuracy and stability. (C) demonstrates the IEE from Node 1 to the other nine nodes, while (D) shows the IEE received by Node 7 from the other nodes. IEE accurately captures the interventional flow and IntDC in cascade topology.
IEE can quantify IntDC effectively without requiring interventions or perturbations
IEE is specifically designed for calculating IntDC solely from observed time-series data without requiring additional perturbations to the system. To validate its accuracy, we conducted a comparison against true perturbed deviations using chaotic neural networks (chNNs) as a model system (?, ?).
The chNN comprises an output variable $\bm{x}=(x_{1},x_{2},\dots,x_{N_{\text{node}}})^{T}$ and two internal variables, with $N_{\text{node}}$ representing the number of neurons in the network. We chose $N_{\text{node}}=10$ in simulations. Details on the dynamics can be found in SM. We denoted the observed time-series data from the stationary system without intervention as $\bm{x}_{\text{obs}}$. When the neuron $i$ was removed (set as a constant zero), the perturbed data were denoted as $\bm{x}_{\text{per}}^{(i)}$. Removing one node was taken as an intervention to the system. The stationary probability density functions of the $j$th neuron before and after the intervention were represented by $p_{j}(\bm{x}_{\text{obs}})$ and $p_{j}(\bm{x}_{\text{per}}^{(i)})$, respectively. We used the Kullback-Leibler divergence (KLD), i.e.
$D_{ij}$ $\triangleq\text{KLD}[p_{j}(\bm{x}_{\text{obs}})||p_{j}(\bm{x}_{\text{per}}^{(i)})]$
$=\iint p_{j}(\bm{x}_{\text{obs}})\log\frac{p_{j}(\bm{x}_{\text{obs}})}{p_{j}(\bm{x}_{\text{per}}^{(i)})}\mathop{}\!\mathrm{d}\bm{x}_{\text{obs}}\mathop{}\!\mathrm{d}\bm{x}_{\text{per}}^{(i)},$ (9)
to quantify the true influence of the intervention, i.e. IntDC, from $i$ to $j$. By conducting $100$ randomly simulated chNNs with interventions (10 removed nodes for each chNN), we recorded $1000$ KLD values and compared them with the results of IEE (calculated from $\bm{x}_{\text{obs}}$ solely). In Fig. 4(A), gray dots are the $1000$ samples of causal edges between neurons, while the red line stands for the linear regression. We observed that $\text{IEE}[x_{i}\rightarrow x_{j}]$ exhibited a positive linear correlation with the KLD $D_{ij}$ (with $R^{2}=0.869$). Results for GC (with $R^{2}=0.572$), TE (with $R^{2}=0.912$), and CCM (with $R^{2}=0.622$) are shown in Figs. 4(B-D), which imply that GC and CCM cannot linearly reflect KLD. Furthermore, the violin plots of cosine similarity $S_{c}$ (see SM) between KLD and the four causal indices on $100$ chNNs are displayed in Fig. 4(E). IEE ($0.904\pm 0.024$) had the highest mean similarity with KLD, compared to GC ($0.711\pm 0.076$), TE ($0.888\pm 0.022$), and CCM ($0.687\pm 0.033$). These results indicate that IEE can accurately quantify IntDC from $\bm{x}_{\text{obs}}$ alone, alleviating the need for additional interventions to the dynamics.
Figure 4: The performance of IEE when measuring the influence of perturbations on the chaotic neural networks (chNNs). The Kullback-Leibler divergence (KLD) is utilized to measure the true influence or IntDC between neurons. In (A-D) , $1000$ samples for different perturbations on chNNs are represented by gray dots, along with the linear regressions (the red lines) between KLD and IEE/GC/TE/CCM, respectively. (E) displays the violin plots of cosine similarity between KLD and IEE/GC/TE/CCM. The mean values and standard variations of the cosine similarity are IEE ($0.904\pm 0.024$), GC ($0.711\pm 0.076$), TE ($0.888\pm 0.022$), and CCM ($0.687\pm 0.033$). IEE can linearly reflect the KLD with high similarity, indicating its effectiveness in accurately quantifying IntDC without additional perturbed data on chNNs.
Application of IEE in inferring neural connectomes of C. elegans
We applied the IEE criterion to infer the neural connectomes of Caenorhabditis elegans (C. elegans), a model organism known for its comprehensively studied nervous system  (?). Calcium fluorescence imaging time series data from $31$ neurons with specific functions of a freely moving C. elegans were collected from Kato et al.  (?) (see Figs. 5(A-B) and fig. S7) . The Fig. 5(C) shows the cosine similarity matrix between different neurons. We clustered the neurons into $7$ clusters after data preprocessing (see details in Supplementary Text and table S5). Figure 5(D) presents the ground truth of neural connectomes in a directed graph, which was detected by electron microscopy in references  (?, ?, ?). We calculated the IEE between different clusters to infer the IntDC network. The ROC curve of IEE is shown in Fig. 5(F) with an AUC value $0.882$. The maximum Youden index, maximum concordance probability, and minimum distance to the point $(0,1)$ gave the same optimal operating point (OOP). At the OOP, Fig. 5(E) displays the inferred connectomic network by IEE. Comparisons with other ObsDC indices, i.e. GC/TE/CCM, revealed that IEE exhibited the highest AUC value, best OOP, largest similarity to the ground truth, lowest false positive at the OOP, and lowest false negative at the OOP (see Table 2, figs. S8-S10, and table S6). These findings demonstrate the efficacy of IEE in reconstructing the neural connectomes of C. elegans.
Figure 5: Application of the IEE on inferring the neural connectomes of C. elegans. (A) is an illustration of C. elegans. (B) displays the calcium imaging time series of $31$ key neurons, with their names listed on the left and colors representing different clusters. (C) is the cosine similarity matrix between neurons, aiding the clustering process. (D) is the true connectomes between $7$ clusters, while (E) is the inferred causal network at the the optimal operating point (OOP) determined by IEE. The red edge represents a false positive, and the black dashed edges represent false negatives. (F) shows the ROC curve (the blue line) for IEE. The same OOP is obtained under three criteria, i.e. the maximum Youden index (the blue circle), the maximum concordance probability (the green dot), and the minimum distance to the point $(0,1)$ (the red cross). The shaded area around the ROC curve represents the $95\%$ confidence interval obtained through bootstrapping. The AUC value of IEE is $0.882$.
Application of IEE to COVID-19 transmission in Japan
We validated that IEE was a promising indicator for assessing the transmission dynamics of infectious diseases. Our study collected the daily confirmed COVID-19 cases from all $47$ prefectures in Japan, spanning a duration of 1209 days from January 16, 2020 to May 8, 2023 (fig. S11).
First, we calculated the IEE indices for IntDC among different prefectures. IEE effectively ranked the influence across prefectures, with the top five affected areas by Tokyo being Kanagawa, Chiba, Saitama, Aichi, and Osaka. Remarkably, this ranking coincides with the geographical proximity and socio-economic connections to Tokyo. Kanagawa, Chiba and Saitama are in the same metropolitan area with Tokyo, while Aichi and Osaka are far from Tokyo but connected to Tokyo by expressway and Shinkansen.
Then, IEE was validated to coincide with the effective distance matrix $D^{\text{COVID}}$ between prefectures. The $D^{\text{COVID}}\in\mathbb{R}^{47\times 47}$ was non-symmetric and designed by incorporating factors such as geometric distances, human mobility, population sizes, and infectious rates across prefectures (based on the gravity model  (?, ?), see Supplementary Text and fig. S11 for details). The element $D_{ij}^{\text{COVID}}$ served as a benchmark for quantifying COVID-19 transmission from prefecture $i$ to $j$. The IEE values demonstrated a strong linear correlation with $\ln D^{\text{COVID}}$, particularly when assessing the causality from Tokyo/Osaka to other prefectures (Figs. 6(A-B)). This correlation remained significant for other regions such as Aichi, Hokkaido, Fukuoka, and Okinawa (see fig. S12).
Further, IEE outperformed ObsDC indices in practicality. Notably, the absolute value of the Pearson correlation coefficient (PCC) between IEE and $\ln D^{\text{COVID}}$ was $0.724$ substantially higher than those obtained by GC $0.161$, TE $0.284$, and CCM $0.568$ (average over specific prefecture, see the second to last column in Table 3 and the boxplot in Fig. 6(C)). Table 3 presented more detailed comparisons to support the effectiveness of IEE (with Tokyo for specific illustration in fig. S13).
These results underscore the reliability and efficacy of IEE in providing quantitative insights into COVID-19 transmission dynamics in Japan. Importantly, IEE performs as a simple and valuable tool for causal analyses solely from daily confirmed time-series data, without the need for complex models considering various factors such as geodesic distances, human mobility, population sizes, and infectious rates.
Figure 6: Application of the IEE on the COVID-19 transmission in Japan and the gene regulatory network (GRN) of circadian rhythms. (A) shows that IEE values from Tokyo to other prefectures are highly linear-correlated with $\ln D^{\text{COVID}}$, where $D^{\text{COVID}}$ is the effective distance obtained from the gravity model in consideration of various factors. (B) is the result for IEE values from Osaka to the other prefectures. (C) is the boxplot for $|\text{PCC}(c_{ij}^{\text{Inf}},\ln D_{ij}^{\text{COVID}})|$ with average over fixed $i$, where PCC is the Pearson correlation coefficient and $c^{\text{Inf}}$ is the inferred causal strengths by IEE, GC, TE, or CCM. The PCC linearity of IEE ($0.724\pm 0.116$) is significantly higher than that of GC ($0.161\pm 0.099$), TE($0.284\pm 0.168$), and CCM ($0.568\pm 0.174$). (D) shows the GRN surrounding Clock, a key circadian gene. (E) demonstrates the ROC curves and AUC values of the IntDC index IEE, and three ObsDC indices GC/TE/CCM.
Application of IEE to investigating circadian rhythms
We investigated the gene regulatory networks (GRNs) related to key genes on the circadian rhythm. The time series of gene expressions were measured by microarray from cultured rat cells  (?, ?). Through decades of molecular and genetic studies, many key circadian genes, such as Clock, Bmal1(Arntl), Dec1(Bhlhb2), Dec2(Bhlhb3), Cry1, Cry2, Per1, Per2, Per3, have been identified on mammals  (?, ?). Figure 6(D) displayed the GRN around Clock at the protein level, where the transcription factor Clock is phosphorylated by PFK family genes. The ROC curves and AUC values for IEE/GC/TE/CCM were showed in Fig. 6(E). IEE designed for quantifying IntDC presented a higher AUC value ($0.737$) than ObsDC indices ($0.490$ for GC, $0.610$ for TE, and $0.644$ for CCM). In fig. S14, we also tested the indices on the GRN surrounding Cry1/Cry2, where IEE outperformed the others.
Concluding remarks and perspective
In summary, we have established a framework of IntDC based on dynamical systems theory and introduced the IEE criterion to quantify IntDC in this study. IEE measures the information flow between interventional causes and effects in the delay-embedding space, making it suitable for analyzing non-linear and non-separable systems (?, ?). Moreover, IEE is able to infer IntDC solely from observational time-series data, without requiring additional perturbations to the system. This property makes IEE particularly suitable for non-intervention systems. Both theoretical derivations and numerical examples presented in this study provide strong evidence to support the high effectiveness, accuracy and robustness of IEE in detecting IntDC and reconstructing networks. Furthermore, IEE serves as an effective tool for evaluating and ranking the causal dependence between variables within a dynamical system. Through real-world examples, we illustrated the promising applicability of IEE in diverse fields such as regulatory inference and disease transmission studies.
There are several compelling issues that warrant further investigation and future development.
(1) The theory of IntDC and the IEE criterion can be extended to analyze causal relations between groups of variables. Specifically, in Eq. (6), the variables $x$ and $y$ can be generalized to high-dimensional vectors, enabling the investigation of complex interactions among multiple variables.
(2) The data in computing IEE can be non-uniformly sampled in the temporal space. We used the time-delayed embedding Eq. (2) and Eq. (3) in this paper, while according to the general Takens’ embedding theorem, any $L+1$ observations of $x_{t}$ and $y_{t}$ can form a topological diffeomorphism to $\mathcal{M}_{X}$ and $\mathcal{M}_{Y}$, respectively. Equation (6) shows that only neighbors in the embedding space matter. Thus, IEE algorithm can be extended to infer causality using diverse datasets sampled at different intervals.
(3) To distinguish direct and indirect causality, we can generalize IEE to its conditional version
$\text{cIEE}[x\rightarrow y|z]=\text{CMI}(\delta\bm{X}_{t},\delta\bm{Y}_{t+1}|\bm{Y}_{t+1},\bm{Z}_{t},\delta\bm{Z}_{t}),$ (10)
where “cIEE” is short for the conditional IEE, $z$ is a third variable whose time-delay vector is $\bm{Z}_{t}$, and $\delta\bm{Z}_{t}$ represents the intervened deviation on $\bm{Z}_{t}$. Equation (10) measures the direct IntDC from $x$ to $y$ conditioned on $z$. Further, together with the well-known PC algorithm  (?), cIEE is one feasible way to remove high-order indirect causal edges iteratively and get the direct causal network.
(4) Moreover, causal inference with deep learning technology represents a frontier in research. Notable works include CausalEGM  (?) and Causalformer (?). Combining IntDC and neural networks is a promising field. Causality theory brings interpretability for artificial intelligence, while big model can provide new discoveries in causal relationships. We are now entering a brilliant “big causality era”.
Acknowledgments
Funding:
This work was supported by the National Key R&D Program of China 2022YFC2704604, the National Natural Science Foundation of China No. 12301620 and No. 32100509, the Moonshot R&D Grant Number JPMJMS2021, Institute of AI and Beyond of UTokyo, the International Research Center for Neurointelligence (WPI-IRCN) at The University of Tokyo Institutes for Advanced Study (UTIAS), JSPS KAKENHI Grant Number JP20H05921, the Cross-ministerial Strategic Innovation Promotion Program (SIP), and the 3rd period of SIP “Smart energy management system” Grant Number JPJ012207.
Author contributions:
Conceptualization: J.S., K.A., L.C., and W.L.
Methodology: J.S., K.A., L.C., and W.L.
Investigation: J.S., Y.L., J.Z., and S.L.
Visualization: J.S., Y.L., J.Z., S.L.
Supervision: K.A., L.C., and W.L.
Writing – original draft: J.S, Y.L., J.Z., and L.C.
Writing – review & editing: All authors.
Competing interests:
The authors declare no competing interest.
Data and materials availability:
The code for IEE is available at https://github.com/smsxiaomayi/IEE.
Supplementary materials
Materials and Methods
Supplementary Text
Figs. S1 to S14
Tables S1 to S6

### v2

@@
-Deciphering interventional dynamical causality from non-intervention systems
+Deciphering Interventional Dynamical Causality from Non-intervention Complex Systems
 […]
 -3pt plus 3pt minus 2pt
-Deciphering interventional dynamical causality from non-intervention systems
-Jifan Shi1,2,3, Yang Li4, Juan Zhao5, Siyang Leng1,6, Kazuyuki Aihara4∗, Luonan Chen4,7,8∗, Wei Lin1,2,3∗
-1Research Institute of Intelligent Complex Systems, Fudan University, Shanghai 200433, China
-2Shanghai Artificial Intelligence Laboratory, Shanghai 200232, China
-3State Key Laboratory of Medical Neurobiology and MOE Frontiers Center for Brain Science, Fudan University, Shanghai 200032, China
-4International Research Center for Neurointelligence, The University of Tokyo Institutes for Advanced Study, The University of Tokyo,Tokyo 113-0033, Japan
-5School of Pharmacy, Shanghai University of Traditional Chinese Medicine, Shanghai 201203, China
-6Academy for Engineering and Technology, Fudan University, Shanghai 200433, China
-7Key Laboratory of Systems Biology, Shanghai Institute of Biochemistry and Cell Biology, Center for Excellence in Molecular Cell Science, Chinese Academy of Sciences, Shanghai 200031, China
-8Key Laboratory of Systems Health Science of Zhejiang Province, School of Life Science, Hangzhou Institute for Advanced Study, University of Chinese Academy of Sciences, Chinese Academy of Sciences, Hangzhou 310024, China
-To whom correspondence should be addressed; E-mail: kaihara@g.ecc.u-tokyo.ac.jp or lnchen@sibs.ac.cn or wlin@fudan.edu.cn.
+Deciphering Interventional Dynamical Causality from Non-intervention Complex Systems
+Jifan Shi1∗, Yang Li2, Juan Zhao3, Siyang Leng1,4, Rui Bao5, Kazuyuki Aihara2∗, Luonan Chen6,7∗, Wei Lin1∗
+1Research Institute of Intelligent Complex Systems & CISOR, Fudan University, Shanghai 200433, China
+2International Research Center for Neurointelligence, The University of Tokyo Institutes for Advanced Study, The University of Tokyo, Tokyo 113-0033, Japan
+3School of Pharmacy, Shanghai University of Traditional Chinese Medicine, Shanghai 201203, China
+4Institute of AI and Robotics, College of Intelligent Robotics and Advanced Manufacturing, Fudan University, Shanghai 200433, China
+5Frontiers Science Center for Deep Ocean Multispheres and Earth System, Key Laboratory of Marine Chemistry Theory and Technology, Ministry of Education, Ocean University of China, Qingdao 266100, China
+6School of Mathematical Sciences and School of AI, Shanghai Jiao Tong University, Shanghai 200240, China
+7Key Laboratory of Systems Health Science of Zhejiang Province, Hangzhou Institute for Advanced Study, University of Chinese Academy of Sciences, Chinese Academy of Sciences, Hangzhou 310024, China
+To whom correspondence should be addressed. E-mail: jfshi@fudan.edu.cn; kaihara@g.ecc.u-tokyo.ac.jp; lnchen@sibs.ac.cn; wlin@fudan.edu.cn.
 Detecting and quantifying causality is a focal topic in the fields of science, engineering, and interdisciplinary studies.
 However, causal studies on non-intervention systems attract much attention but remain extremely challenging.
-To address this challenge, we propose a framework named Interventional Dynamical Causality (IntDC) for such non-intervention systems, along with its computational criterion, Interventional Embedding Entropy (IEE), to quantify causality.
-The IEE criterion theoretically and numerically enables the deciphering of IntDC solely from observational (non-interventional) time-series data, without requiring any knowledge of dynamical models or real interventions in the considered system.
-Demonstrations of performance showed the accuracy and robustness of IEE on benchmark simulated systems as well as real-world systems, including the neural connectomes of C.
-elegans, COVID-19 transmission networks in Japan, and regulatory networks surrounding key circadian genes.
-Introduction
+Delay-embedding technique provides a promising approach.
+In this study, we propose a framework named Interventional Dynamical Causality (IntDC) in contrast to the traditional Constructive Dynamical Causality (ConDC).
+ConDC, including Granger causality, transfer entropy and convergence of cross-mapping, measures the causality by constructing a dynamical model without considering interventions.
+A computational criterion, Interventional Embedding Entropy (IEE), is proposed to measure causal strengths in an interventional manner.
+IEE is an intervened causal information flow but in the delay-embedding space.
+Further, the IEE theoretically and numerically enables the deciphering of IntDC solely from observational (non-interventional) time-series data, without requiring any knowledge of dynamical models or real interventions in the considered system.
+In particular, IEE can be applied to rank causal effects according to their importance and construct causal networks from data.
+We conducted numerical experiments on Logistic dynamics, coupled-Henon maps, and chaotic neural networks to demonstrate that IEE can find causal edges accurately, eliminate effects of confounding, and quantify causal strength robustly over traditional indices.
+We also applied IEE to real-world tasks, including estimating neural connectomes of C.
+elegans, detecting COVID-19 transmission networks in Japan, and investigating regulatory networks surrounding key circadian genes.
+IEE performed as an accurate and robust tool for causal analyses solely from the observational data.
+The IntDC framework and IEE algorithm provide an efficient approach to the study of causality from time series in diverse non-intervention complex systems.
+1 Introduction
 Directional or indirectional interactions among different components give rise to a variety of complex phenomena in nature and social society.
@@
 Modeling, detecting, and quantifying causality from observational data are crucial for describing, interpreting, predicting, and even controlling complex systems.
-Statisticians believe that the causality is contained and obtainable in universal random variables regardless of the time label.
-The Neyman-Rubin’s potential outcome framework  (?, ?, ?), the Wright’s structural equation model  (?, ?) and the Pearl’s causal diagram model  (?, ?) are the most famous statistical approaches, which have been proved to be of mathematical equivalence  (?).
-Especially, the instrumental variable method, that was awarded the 2021 Nobel Prize in economic sciences, is widely applied in the study of causality among economics, environments, and other disciplines  (?).
+Statisticians believe that causality is contained and obtainable in universal random variables regardless of the time label.
+Neyman-Rubin’s potential outcome framework  (?, ?, ?), the Wright’s structural equation model  (?, ?) and the Pearl’s causal diagram model  (?, ?) are the most famous statistical approaches, which have been proved to be of mathematical equivalence  (?).
+Especially, the instrumental variable method is widely applied in the study of causality among economics, environments, and other disciplines  (?).
 Statistical methods seek to discern binary causal relations among random variables on a directed acyclic graph, without relying on temporal data.
@@
 Transfer entropy (TE)  (?) generalizes GC to the nonlinear case by quantifying the prediction uncertainty through Shannon entropy.
-Neither GC nor TE addresses the “non-separability” problem, which means that removing the causal variable from the system inevitably influences the dynamics of downstream variables  (?, ?, ?, ?).
+Neither GC nor TE addresses the “non-separability” problem, which means that removing the causal variable from the system inevitably influences the dynamics of downstream variables  (?, ?, ?, ?, ?, ?).
 To measure causality in non-separable systems, numerous approaches have emerged in the last decade within the framework of delay embedding.
 These include convergence of cross mapping (CCM)  (?, ?), partial cross mapping (PCM)  (?), continuity scaling  (?), inverse continuity  (?), topological expansion  (?), joint distance distribution  (?), embedding entropy (EE)  (?) and other indices  (?, ?).
-Nonetheless, most indices designed for quantifying the dynamical causality primarily estimate directional causal relations at the observational level, i.e.
-observational dynamical causality (ObsDC).
-Conventionally, to detect and quantify causality at the interventional level, which is noted as the interventional dynamical causality (IntDC) in this study, intelligent modulation and manipulation of the dynamical system should be taken into account.
+Besides, Runge et al.
+designed the PCMCI method to further remove “common drivers” by selecting parent nodes iteratively (?).
+Friston et al.
+proposed the dynamical causal modeling (DCM), which is a Bayesian fitting from data with pre-selected models  (?).
+DCM was originally developed for modeling neural dynamics  (?).
+Krakovská et al.
+conducted a comparative analysis of six methods for detecting causality in bivariate systems (?).
+Nonetheless, most indices designed for quantifying causality from time series primarily focus on estimating directional causal relationships at the constructive level, referred to as constructive dynamical causality (ConDC).
+Traditionally, to detect and quantify causality at the interventional level, which is termed interventional dynamical causality (IntDC) in this study, requires intelligent modulation and manipulation of the dynamical system.
 By allowing external intervention to the system and recording data under different perturbations, frameworks such as perturbation cascade inference (PCI)  (?) and dynamical causal effect (DCE)  (?, ?) provide relevant computational schemes.
+To detect asymmetry information transfer in known two-dimensional dynamical systems, Liang and Kleeman proposed an analytic approach named Liang-Kleeman information flow  (?, ?, ?), which serves as a prototype for IntDC.
+This method quantifies causality by freezing one variable as a parameter, and evaluating the resulting outcomes using the Frobenius-Perron operator.
+The Liang information flow has been further extended to multivariate, stochastic, and quantum systems  (?, ?, ?, ?).
 However, due to ethical or practical limitations, many real systems should be analyzed without any external intervention.
@@
 Numerical IEE does not require specific prior knowledge of dynamics, and additional perturbation to the system is also unnecessary.
-Compared to ObsDC indices, IEE designed for IntDC has the capability to rank the importance of causal effects and construct directional causal networks more effectively.
+Compared to ConDC indices, such as GC, TE, and CCM, IEE designed for IntDC has the capability to rank the importance of causal effects and construct directional causal networks more effectively.
 We demonstrate numerical experiments on both simulated examples and real datasets, including estimating neural connectomes of C.
 elegans, evaluating COVID-19 transmission in Japan, and constructing regulatory networks surrounding key circadian genes.
-Methods
-Observational dynamical causality in delay-embedding space
-We consider a discrete two-variable dynamical system as
+2 Methods
+2.1 Constructive dynamical causality
+The definition of ConDC for a complex system in the original time-series space is given as follows:
+Definition 1 (Constructive dynamical causality, ConDC  (?)).
+For a complex dynamical system
+$\bm{x}_{t+1}=\bm{f}(\bm{x}_{t},\bm{x}_{t-1},\dots,\bm{x}_{t-p})+\bm{\varepsilon}_{t},$ (1)
+where $\bm{x}=(x^{(1)},x^{(2)},\dots,x^{(n)})^{T}$ represents the system with $n$ components, $\bm{f}=(f_{1},f_{2},\dots,f_{n})^{T}$ is a vector function, $p$ is referred to the memory time, and $\bm{\varepsilon}_{t}$ is an independent noise term, there exists ConDC from component $x^{(j)}$ to $x^{(i)}$, if $\exists k\in\{1,2,\dots,p\}$ such that $\partial f_{i}/\partial x^{(j)}_{t-k}\neq 0$ for almost any $t$.
+For simplicity, we consider a discrete two-variable dynamical system as an example:
 $\begin{cases}x_{t+1}=g(x_{t},x_{t-1},\dots,x_{t-p},\varepsilon_{x,t}),\\ y_{t+1}=f(x_{t},x_{t-1},\dots,x_{t-p},y_{t},y_{t-1},\dots,y_{t-p},\varepsilon_{y,t}),\end{cases}$ (2)
 where $x,y$ are two variables, $p$ denotes the time step during which causality is considered, and $\varepsilon_{\cdot,t}$ stand for small noise terms.
-According to Shi et al.
-(?), there exists ObsDC from $x$ to $y$ since the evolving equation of $y_{t+1}$ depends on the historical behavior of $x$, while there is no ObsDC from $y$ to $x$ as the dynamics of $x_{t+1}$ is independent of $y$.
-GC and TE measure the predictability of $f$ in the original time-series space to quantify the ObsDC from $x$ to $y$ through linear regression and entropy uncertainty, respectively (Fig.
+According to Definition 1, there exists ConDC from $x$ to $y$ since the evolving equation of $y_{t+1}$ depends on the historical behavior of $x$, while there is no ConDC from $y$ to $x$ as the dynamics of $x_{t+1}$ is independent of $y$.
+GC and TE measure the predictability of $f$ in the original time-series space to quantify the ConDC from $x$ to $y$ through linear regression and entropy uncertainty, respectively (Fig.
 1(B)).
@@
 However, in general coupled systems, the separability is not satisfied, as removing one variable changes the values of affected ones  (?, ?, ?).
-To address the challenge of detecting ObsDC in universal non-separable systems, various approaches from the delay-embedding space have been proposed.
+To address the challenge of detecting ConDC in universal non-separable systems, various approaches from the delay-embedding space have been proposed.
 In the delay-embedding framework, the bivariate system Eq.
 (2) is assumed to evolve into an attractive manifold with an inner dimension $d$.
-Let the time-delay vectors of $x$ and $y$ be
+Let the time-delayed vectors of $x$ and $y$ be
 $\bm{X}_{t}=(x_{t},x_{t-1},\dots,x_{t-L})^{T}\in\mathcal{M}_{X}\subseteq\mathbb{R}^{L+1},$ (3)
@@
 $\bm{Y}_{t+1}=(y_{t+1},y_{t},y_{t-1},\dots,y_{t-L+1})^{T}\in\mathcal{M}_{Y}\subseteq\mathbb{R}^{L+1},$ (4)
-where $L$ is the time-delay length, $\mathcal{M}_{X}$ and $\mathcal{M}_{Y}$ represent the manifolds formed by $\bm{X}_{t}$ and $\bm{Y}_{t}$, respectively.
-According to the seminal stochastic version of Takens’ embedding theorem  (?, ?, ?, ?), for delay dimension $L\geqslant 2d$, we can obtain
+where $L$ is the time-delayed length, $\mathcal{M}_{X}$ and $\mathcal{M}_{Y}$ represent the manifolds formed by $\bm{X}_{t}$ and $\bm{Y}_{t}$, respectively.
+According to the seminal stochastic version of Takens’ embedding theorem  (?, ?, ?, ?), we can obtain the following theorem:
+Theorem 1 (ConDC in delay-embedding space).
+If $x$ is the ConDC of $y$ in dynamics Eq.
+(2), and $\bm{X}_{t}$, $\bm{Y}_{t}$ are the time-delayed vectors, respectively, then there exists a smooth projection operator $\bm{F}$ in generic sense such that
 $\bm{X}_{t}=\bm{F}(\bm{Y}_{t+1}),$ (5)
-where $\bm{F}$ is a smooth projection operator.
+when the time-delayed length satisfies $L\geqslant 2d$, where $d$ is the inner dimension of the attractive manifold.
 The detailed derivation can be referred to the Supplementary Text.
+The “generic sense” means that a smooth projection exists for a dense and open set of all possible time-delayed ways (?, ?).
 Equation (5) shows that the causal variable $\bm{X}_{t}\in\mathcal{M}_{X}$ can be reconstructed by the effect variable $\bm{Y}_{t+1}\in\mathcal{M}_{Y}$.
 However, as $\bm{F}$ is not reversible in generic sense, we can not determine $\bm{Y}_{t+1}$ only by the information from $\bm{X}_{t}$.
-The causal dependence in $f$ between $x$ and $y$ is transformed into the reconstructability of $\bm{F}$ (Fig.
+According to Theorem 1, the causal dependence in $f$ between $x$ and $y$ is transformed into the reconstructability of $\bm{F}$ (Fig.
 1(D)).
+Theorem1, established for the dynamical system in Eq.
+(2), can be generalized to multi-variable systems as described in Eq.
+(1), allowing for feedback interactions between different variables.
 Instead of fitting a model by removing the causal variable as in GC and TE, detecting the existence and quantifying the continuity characteristics of $\bm{F}$ in the delay-embedding space can be adequate for causal identification, especially for universal non-separable systems.
 Related algorithms include CCM, PCM, and EE.
-Figures 1(A), (B) and (D) summarize dynamical causality at the observational level, where the ObsDC from variable $x$ to $y$ is estimated only from the historical observed time series.
+Figures 1(A), (B) and (D) summarize dynamical causality at the constructive level, where the ConDC from variable $x$ to $y$ is estimated only from the historical observed time series.
 But fitting Eq.
@@
 (5) from data is criticized as mere association or prediction  (?, ?).
-Figure 1: Illustration for the observational dynamical causality (ObsDC) and interventional dynamical causality (IntDC).
+Figure 1: Illustration for the constructive dynamical causality (ConDC) and interventional dynamical causality (IntDC).
 (A) An illustrative example of a complex system where $(x_{t},y_{t})$ are observed time series.
-The $z_{t}$ stands for the unobserved variable.
-(B) ObsDC from $x$ to $y$ indicates that the dynamics of $y$ depends on $x$.
-Algorithms such as Granger causality (GC) and transfer entropy (TE) infer the ObsDC in the original time-series space.
+The $z_{t}$ stands for the other variable that is not the primary focus of analysis.
+(B) ConDC from $x$ to $y$ indicates that the dynamics of $y$ depends on $x$.
+Algorithms such as Granger causality (GC) and transfer entropy (TE) infer the ConDC in the original time-series space.
 (C) IntDC from $x$ to $y$ indicates that the deviation of $y$, i.e.
@@
 $\delta x$.
-Usually additional data from a perturbed system $(\tilde{x}_{t},\tilde{y}_{t})$ is necessary for detecting IntDC.
-(D) ObsDC can be detected in the embedding space by cross mapping $\mathbb{F}$.
-The causal vector $X_{t}$ can be reconstructed by the effect vector $\bm{Y}_{t+1}$.
+Usually additional data from a perturbed system $(\tilde{x}_{t},\tilde{y}_{t})$ is necessary for detecting IntDC in the original time-series space.
+The parameter $p$ is the time lag after intervention occurs.
+(D) ConDC can be detected in the embedding space by cross mapping $\mathbb{F}$.
+The causal vector $\bm{X}_{t}$ can be reconstructed by the effect vector $\bm{Y}_{t+1}$.
+Parameter $L$ is time delay used in the delay-embedding.
 (E) IntDC can be modeled in the embedding space.
 The interventional causal vector $\delta\bm{X}_{t}$ should be reconstructable by the effect vector $\delta\bm{Y}_{t+1}$ around $\bm{Y}_{t+1}$.
-The $\delta\bm{Y}_{t+1}$ is approximated by the distance from $\bm{Y}_{t+1}$ to its neighbors $\bm{Y}_{t_{k}+1}$, while $\delta\bm{X}_{t}$ is obtained by $X_{t_{k}}$ with the same time label $t_{k}$.
+In numerical computation, $\delta\bm{Y}_{t+1}$ is approximated by the distance from $\bm{Y}_{t+1}$ to its neighbors $\bm{Y}_{t_{k}+1}$, while $\delta\bm{X}_{t}$ is determined using $\bm{X}_{t_{k}}$ corresponding to the same time index $t_{k}$.
 (F) Interventional embedding entropy (IEE) measures the average information retention contained in $\delta\bm{Y}_{t+1}$ from $\delta\bm{X}_{t}$, quantifying the IntDC from variable $x$ to $y$.
 Notably, IEE does not require additional perturbations to the system and can be inferred from observational data solely.
-Interventional dynamical causality in delay-embedding space
+2.2 Interventional dynamical causality
 The detection of hidden essential-level causality necessitates intervention or manipulation to the system  (?, ?, ?, ?).
-We define that there is an IntDC from $x$ to $y$ during the time $[t-p,t]$ in the original time-series space, if $\delta y_{t}=\tilde{y}_{t}-y_{t}$ depends on $\delta x_{t}=\tilde{x}_{t}-x_{t}$, where the dynamics of $x_{t}$ is perturbed to $\tilde{x}_{t}$, and $\tilde{y}_{t}$ is the dynamics of affected $y$ after the perturbation (Fig.
-1(C)).
-The intervention $\delta x_{t}$ can be either a pulse stimulation or persistent disturbances to the system.
+In this study, we propose the definition of IntDC for a complex system in the original time-series space as the following:
+Definition 2 (Interventional dynamical causality, IntDC).
+For a complex dynamical system Eq.
+(1), one component $x^{(j)}$ is the IntDC of another component $x^{(i)}$, if $\exists k\in\{1,2,\dots,p\}$ such that $\delta x^{(i)}_{t+1}=\tilde{x}^{(i)}_{t+1}-x^{(i)}_{t+1}$ depends on $\delta x^{(j)}_{t-k}=\tilde{x}^{(j)}_{t-k}-x^{(j)}_{t-k}$ for almost any $t$, where $\tilde{x}^{(j)}_{t-k}$ is the intervened dynamics of $x^{(j)}$, and $\tilde{x}^{(i)}_{t+1}$ is the dynamics of affected $x^{(i)}$ after the intervention on $x^{(j)}$.
+The intervention $\delta x^{(j)}_{t-k}$ can be either a pulse stimulation or persistent disturbances to the system.
+Fig.
+1(C) illustrates the IntDC in the time-series space for the simplified bivariate system Eq.
+(2), involving variables $x$ and $y$.
 To measure the IntDC, data before and after the intervention, i.e.
 $(x_{t},y_{t})$ and $(\tilde{x}_{t},\tilde{y}_{t})$, are typically required, as in PCI and DCE.
+PCI measures the distances of every node in the network from each perturbed node to reconstruct the directed causal network  (?).
+DCE classified causal coupling into four situations based on the type of effects (stationary statistic change, phase orbit change) and the type of interventions (state space interventions, parametric interventions)  (?).
 These conventional algorithms rely on reproducible dynamics or data availability under different settings.
-To address challenges of causal inference in widespread non-intervention, non-linear and non-separable systems, we turn to the study of intervened dynamics Eq.
+To address challenges of causal inference in widespread non-intervention, non-linear and non-separable systems, we turn to the study of intervened dynamics of Eq.
 (5) in the delay-embedding space.
-Denote $\delta\bm{X}_{t}=\widetilde{\bm{X}}_{t}-\bm{X}_{t}$ and $\delta\bm{Y}_{t+1}=\widetilde{\bm{Y}}_{t+1}-\bm{Y}_{t+1}$, where $\widetilde{\bm{X}}_{t}$ and $\widetilde{\bm{Y}}_{t+1}$ are the time-delay vectors of $x$ and $y$ after the intervention, respectively.
-Under infinitesimal intervention, Eq.
-(4) leads to
-$\delta\bm{X}_{t}=\nabla_{\bm{Y}}\bm{F}(\bm{Y}_{t+1})\cdot\delta\bm{Y}_{t+1}.$ (5)
-Equation (5) indicates that $\delta\bm{Y}_{t+1}$ contains complete information of $\delta\bm{X}_{t}$ in the neighborhood of $\bm{Y}_{t+1}$ or reconstructs $\delta\bm{X}_{t}$, if there exists IntDC from the causal variable $x$ to the effect variable $y$ (Fig.
+Denote $\delta\bm{X}_{t}=\widetilde{\bm{X}}_{t}-\bm{X}_{t}$ and $\delta\bm{Y}_{t+1}=\widetilde{\bm{Y}}_{t+1}-\bm{Y}_{t+1}$, where $\widetilde{\bm{X}}_{t}$ and $\widetilde{\bm{Y}}_{t+1}$ are the time-delayed vectors of $x$ and $y$ after the intervention, respectively.
+We can obtain the following theorem:
+Theorem 2 (IntDC in delay-embedding space).
+If $x$ is the IntDC of $y$ in dynamics Eq.
+(2), and $\bm{X}_{t},\bm{Y}_{t}$ are the time-delayed vectors, respectively, then there exists a smooth projection operator $\bm{F}$ in generic sense such that
+$\delta\bm{X}_{t}=\nabla_{\bm{Y}}\bm{F}(\bm{Y}_{t+1})\cdot\delta\bm{Y}_{t+1},$ (6)
+when the time-delayed length satisfies $L\geq 2d$ and the intervention is sufficiently small, where $d$ is the inner dimension of the attractive manifold.
+Theorem 2 is directly deduced by Theorem 1.
+However, Eq.
+(6) provides insights on the causality under intervention.
+It indicates that $\delta\bm{Y}_{t+1}$ contains complete information of $\delta\bm{X}_{t}$ in the neighborhood of $\bm{Y}_{t+1}$ and can reconstruct $\delta\bm{X}_{t}$, if there exists IntDC from the causal variable $x$ to the effect variable $y$ (Fig.
 1(E)).
+The conclusion in Eq.
+(6) of Theorem 2, derived from the dynamics in Eq.
+(2), remains valid for any pair of components exhibiting IntDC in the multi-variable system of Eq.
+(1), and feedback between different components is permitted.
 Figures 1(A), (C), and (E) give an overview of the IntDC framework.
-Interventional embedding entropy
+2.3 Interventional embedding entropy
 We introduce the IEE criterion, specifically designed for quantifying IntDC solely from the observational data in the delay-embedding space.
+Definition 3 (Interventional embedding entropy, IEE).
 The IEE criterion is proposed to measure the average information retention or reconstructability from the effect $\delta\bm{Y}_{t+1}$ to the causation $\delta\bm{X}_{t}$, formulated as
-$\text{IEE}[x\rightarrow y]=\text{CMI}(\delta\bm{X}_{t},\delta\bm{Y}_{t+1}|\bm{Y}_{t+1}),$ (6)
-where CMI is the conditional mutual information  (?), i.e.
+$\mathrm{IEE}[x\rightarrow y]:=\mathrm{CMI}(\delta\bm{X}_{t},\delta\bm{Y}_{t+1}|\bm{Y}_{t+1}).$ (7)
+In Eq. (7), CMI is the conditional mutual information  (?), i.e.
 $\text{CMI}(\bm{x},\bm{y}|\bm{z})=\iiint p(\bm{x},\bm{y},\bm{z})\log\frac{p(\bm{x},\bm{y}|\bm{z})}{p(\bm{x}|\bm{z})p(\bm{y}|\bm{z})}\mathop{}\!\mathrm{d}\bm{x}\mathop{}\!\mathrm{d}\bm{y}\mathop{}\!\mathrm{d}\bm{z}.$ (8)
-The $\delta\bm{Y}_{t+1}$ is estimated by the distance from $\bm{Y}_{t+1}$ to its neighbored points $\bm{Y}_{t_{k}+1}$ in the embedding space, and $\delta\bm{X}_{t}$ is calculated by $\bm{X}_{t_{k}}$ with the same time label $t_{k}$ (Figs.
+IEE provides is a quantitative measure of causal strength that is comparable across different edges within the same dynamical system.
+When dealing with data, the $\delta\bm{Y}_{t+1}$ is estimated by the distance from $\bm{Y}_{t+1}$ to its neighbored points $\bm{Y}_{t_{k}+1}$ in the embedding space, and $\delta\bm{X}_{t}$ is calculated by $\bm{X}_{t_{k}}$ with the same time label $t_{k}$ (Figs.
 1(E) and (F)).
@@
 (7) can be numerically approximated only from the observational data, especially for non-intervention systems.
-Table 1 shows the numerical algorithm for IEE, whose detailed derivation can be found in the Supplementary Text.
-Results
-IEE can accurately and robustly quantify causality
+Table 1 presents the numerical algorithm for computing IEE, whose detailed descriptions and computational procedures are provided in the Supplementary Text.
+2.4 Connections with statistical causal models and assumptions of ConDC/IntDC
+The concept of dynamical causality in Eq.
+(1) develops from traditional structural causal models, which typically studies structure equations of the form $Y=f(X,\varepsilon,\theta)$, where $X$ and $Y$ are random variable, $f$ is a function parameterized by $\theta$, and $\varepsilon$ represents noise or residue terms.
+In  (?), the framework of ConDC is systematically studied with three key generalizations beyond classical structural equation modeling: (i) Temporal information is fully considered by dynamical systems, ensuring that causes precede effects in time; (ii) Feedback or causal loops between variables are permitted under the consideration of time delays, relaxing the restriction of directed acyclic graphs (DAGs); (iii) Quantitative measures of causal strength are provided beyond qualitative identification, enabling direct comparisons of causality between different variables.
+The IntDC proposed in the study further integrates the concept of interventions and potential outcomes within the dynamical framework (?).
+IntDC aims to quantify causal strengths at the interventional level directly from observational data (without the need for additional interventions to the system, which are often impractical or infeasible in real-world datasets).
+The following assumptions (A1)-(A8) clarify key connections between dynamical causality and statistical causal inference:
+(A1) Temporal Order assumption: Causal information is embedded in time-series data and causes must precede their effects;
+(A2) Causal Sufficiency (Unconfoundedness) assumption: All common drivers (confounders) are observed such that the causal graph accurately represents relationships among the observed variables;
+(A3) Causal Markov assumption: Each variable $x_{i}$ is independent of its nondescendants given its direct causes (parents);
+(A4) Faithfulness assumption: The causal graph structure can represent the conditional independence contained in the joint probability density;
+(A5) Stationarity assumption: The dynamical system evolves into a stable attractive manifold, allowing stable causal relations to be measured from observational data;
+(A6) Intervention Ignorability assumption: Given the system dynamics Eq.
+(1), initial conditions $\bm{x}_{0}$ and noise $\bm{\varepsilon}_{t}$, the intervention on variable $x^{(j)}$ at time $t-k$ is independent of the potential outcome of $x^{(i)}$ at time $t+1$, i.e.
+$x^{(i)}_{t+1}\text{$\perp\!\!\!\perp$}\delta x^{(j)}_{t-k}|(\bm{f},\bm{x}_{0},\bm{\varepsilon}_{t})$ for $k=1,2,\dots,p$;
+(A7) Stable Infinitesimal Intervention assumption: The average information retention from $x^{(j)}$ to $x^{(i)}$ in dynamics Eq.
+(1) remains stable under sufficiently small interventions $\delta x^{(j)}$;
+(A8) Consistency assumption: The outcome $\tilde{x}^{(i)}_{t+1}$ in Eq.
+(1) can be precisely determined under known dynamics $(\bm{f},\bm{x}_{0},\bm{\varepsilon}_{t})$ and a fixed intervention $\delta x^{(j)}_{t-k}$.
+In practice, (A2) can be relaxed if the primary goal is to quantify the total causal strength (including both direct and indirect causality) between observed variables.
+In the Results section, we demonstrate that the IEE criterion is capable of mitigating the impact of confounding variables, thereby providing robust causal estimates even when some confounders are unobserved.
+Some of these assumptions for causal discovery from observational time-series data have also been discussed by Runge et al.
+( (?, ?)).
+GC, TE, and CCM are all criteria for ConDC, which estimate causal strengths by fitting constructive dynamical models (i.e.
+Eq.
+(2) in the time-series space or Eq.
+(5) in the delay-embedding space).
+GC employs linear vector regression to fit $\bm{f}$ in Eq.
+(2).
+TE extends GC to nonlinear cases by quantifying information transfer between variables.
+CCM constructs a local-linear cross-mapping $\bm{F}$ in Eq.
+(5) by the time-delayed embedding to capture causality.
+For a detailed comparison of these ConDC methods, readers are referred to Table 1 in  (?).
+3 Results
+3.1 IEE captures the numerical behavior of causality in Logistic systems robustly
 To assess the numerical performance of measuring IntDC, we first applied the IEE algorithm on the following two-node Logistic dynamics
@@
 where $\varepsilon_{\cdot,t}$ represents independent Gaussian noises, and parameters $\beta_{xy}$ and $\beta_{yx}$ modulate the IntDC strength from $x$ to $y$ and from $y$ to $x$, respectively (see further details in SM).
-IEE can accurately capture the numerical behavior of dynamical causality in the two-node Logistic system, especially IntDC rather than ObsDC.
-In Fig.
-2(A), we set $\beta_{xy}\equiv 0$ and let $\beta_{yx}$ increase from 00 to $0.3$.
-We observed a monotonic increase in $\text{IEE}[y\rightarrow x]$ (the red line with squares), while $\text{IEE}[x\rightarrow y]$ remained around zero (the blue line with dots).
-In Fig.
-2(B), we set $\beta_{xy}\equiv 0.1$ and let $\beta_{yx}$ increase from 00 to $0.3$.
-$\text{IEE}[y\rightarrow x]$ exhibited a monotonic increase (the red line with squares), while $\text{IEE}[x\rightarrow y]$ stayed above zero (the blue line with dots).
+IEE can capture the numerical behavior of IntDC in the two-node Logistic system.
+When we set $\beta_{xy}\equiv 0$ and let $\beta_{yx}$ increase from 0 to 0.30.30.3.
+we observed a monotonic increase in $\text{IEE}[y\rightarrow x]$ (the red line with squares in Fig.
+2(A)), while $\text{IEE}[x\rightarrow y]$ remained around zero (the blue line with dots in Fig.
+2(A)).
+When we set $\beta_{xy}\equiv 0.1$ and let $\beta_{yx}$ increase from 0 to 0.30.30.3, $\text{IEE}[y\rightarrow x]$ exhibited a monotonic increase (the red line with squares in Fig.
+2(B)), while $\text{IEE}[x\rightarrow y]$ stayed above zero (the blue line with dots in Fig.
+2(B)).
 The gray dashed lines in Figs.
-2(A-B) represent $0.01$ for reference.
-A detailed comparison with GC, TE and CCM can be found in figs.
-S1 and S2, which indicate that GC exhibited non-monotonic behavior, $\text{TE}[y\rightarrow x]$ slightly decreased with a false-negative issue, and $\text{CCM}[x\rightarrow y]$ had a false-positive problem.
-The mean values and standard deviations of the indices under 100 simulations are provided in tables S1-S4.
+2(A-B) represent 0.010.010.01 for reference.
+A detailed comparison with GC, TE and CCM is presented in figs.S1 and S2.
+The mean values and standard deviations of the indices over 100 simulation runs are provided in tables S1-S4.
+Fig.
+S1(b) and fig.
+S2(b) indicate that GC exhibited non-monotonic behavior.
+$\text{TE}[y\rightarrow x]$ demonstrated a decreasing trend when $\beta_{yx}$ increased, specifically for $\beta_{yx}\in(0,0.02]$ in fig.
+S1(c)/table S1 and $\beta_{yx}\in(0,0.03]$ in fig.
+S2(c)/table S3, which may lead to false-negative issues (ignorance of existence causality).
+$\text{CCM}[x\rightarrow y]$ remained significantly different from zero even when $\beta_{xy}=0$ (in the range $\beta_{yx}\in[0.15,0.3]$ in fig.
+S1(d)/table S2).
+$\text{CCM}[y\rightarrow x]$ was also significantly different from zero when $\beta_{yx}=0$ (with $\beta_{xy}=0.1$ in fig.
+S2(d)/table S3).
+These results suggest that CCM may suffer from the false-positive problem (incorrect identification of non-existence causality).
 Within suitable parameter ranges, IEE demonstrates stability and robustness.
-In Figs.
-2(D-G), we validated the robustness of the IEE algorithm under various conditions, including different delayed lags $L$, the numbers of nearest neighbors $K$, lengths of time series $N$, and noise amplitudes $\sigma$.
-The causal strength should exhibit a descending trend with decreasing $\beta_{yx}$ from $0.15$ to $0.05$ through $0.125,0.1,0.075$.
-The experiments support that IEE maintains the correct order under different parameter values with relatively stable variations.
-IEE can eliminate the effects of confounding variables
+We validated the robustness of the IEE algorithm under various conditions, including different delayed lags $L$, the numbers of nearest neighbors $K$, lengths of time series $N$, and noise standard deviations $\sigma$.
+The causal strengths measured by IEE exhibited a consistent descending trend as $\beta_{yx}$ decreased from 0.150.150.15 to 0.050.050.05 through 0.125,0.1,0.0750.125,0.1,0.0750.125 , 0.1 , 0.075, as expected (Figs.
+2(D-G)).
+These results support that IEE reliably preserves the correct ranking of causal strengths across different parameter settings with relatively stable variance.
+3.2 IEE eliminates the influence of confounding variables
 IEE remains applicable and accurate even in presence of confounding variables.
@@
 When $\beta_{xy}=0$, $\text{IEE}[x\rightarrow y]$ was almost zero and significantly different from $\text{IEE}[z\rightarrow x]$ and $\text{IEE}[z\rightarrow y]$ (Fig.
-2(C)); but conventional ObsDC indices such as GC/TE/CCM produced false-positive causality from $x$ to $y$ due to the presence of confounder $z$ (fig.
+2(C)); but conventional ConDC indices such as GC/TE/CCM produced false-positive causality from $x$ to $y$ due to the presence of confounder $z$ (fig.
 S3).
-That is because an intervention on $x$, i.e.
-$\delta x$, will not induce $\delta y$ through $z$ under the IntDC framework.
+According to the IntDC framework, an intervention on $x$ (i.e.
+$\delta x$) will not induce an indirect change in $y$ (i.e.
+$\delta y$) through $z$; in other words, there is no intervention-induced pathway through $x\leftarrow z\rightarrow y$.
+However, traditional ConDC indices by fitting dynamical models may falsely infer an association between $x$ and $y$ due to the influence of the confounder $z$.
 IEE, specially designed for measuring IntDC, remained unaffected by the confounder $z$ and yielded accurate results.
-Figure 2: Performance of IEE on the Logistic systems.
+Figure 2: Performance of IEE on Logistic systems.
 (A) and (B) used the two-node Logistic system.
 The $\beta_{xy}$ and $\beta_{yx}$ control the coupling coefficients from $x$ to $y$ and from $y$ to $x$, respectively.
-(A) shows the mean values of IEE when $\beta_{xy}=0$ and $\beta_{yx}$ ranges from 00 to $0.3$ over 100 simulations.
+(A) shows the mean values of IEE when $\beta_{xy}=0$ and $\beta_{yx}$ ranges from 0 to 0.30.30.3 over 100 simulations.
 $\text{IEE}[y\rightarrow x]$ (the red line with squares) increases monotonically, while $\text{IEE}[x\rightarrow y]$ (the blue line with dots) stays around zero.
 (B) is under $\beta_{xy}=0.1$, in which $\text{IEE}[x\rightarrow y]$ (the blue line with dots) is significantly positive.
-The gray dashed lines in (A) and (B) represent the constant $0.01$ as reference.
+The gray dashed lines in (A) and (B) represent the constant 0.010.010.01 as reference.
 (C) shows the performance of IEE in the three-node Logistic system, where $z$ acts as a confounding variable.
 When $\beta_{xy}=0$, $\text{IEE}[x\rightarrow y]$ is almost zero correctly.
-IEE designed for measuring IntDC is not affected by confounders, which usually leads to false positives.
+IEE designed for measuring IntDC is not affected by confounders, which usually leads to false positives in ConDC.
 The results for $\beta_{xy}=0.5$ is shown for a comparison.
-The gray dashed line is $0.05$ for reference.
-(D-G) demonstrate the robustness of IEE under different delayed lags $L$, the numbers of nearest neighbors $K$, data lengths $N$, and noise amplitudes $\sigma$ in the two-node logistic system, respectively.
-The parameter $\beta_{xy}\equiv 0$, and $\beta_{yx}$ takes values of $0.15$ (red), $0.125$ (blue), $0.1$ (black), $0.075$ (green), and $0.05$ (magenta).
+The gray dashed line is 0.050.050.05 for reference.
+(D-G) demonstrate the robustness of IEE under different delayed lags $L$, the numbers of nearest neighbors $K$, data lengths $N$, and noise standard deviations $\sigma$ in the two-node logistic system, respectively.
+The parameter $\beta_{xy}\equiv 0$, and $\beta_{yx}$ takes values of 0.150.150.15 (red), 0.1250.1250.125 (blue), 0.10.10.1 (black), 0.0750.0750.075 (green), and 0.050.050.05 (magenta).
 Lines denote mean values under 50 simulations, and shaded areas represent the standard deviation.
-IEE maintains correct orders of IntDC under different situations with relatively stable variations.
-IEE can reconstruct causal networks with quantified causal strength
+IEE consistently preserves the correct ranking of IntDC across various settings, with relatively stable variance.
+3.3 IEE enables causal network reconstruction with quantification of causal strengths
 IEE has the capability to reconstruct the causal network structure and rank the causal influence between nodes in complex networks at the interventional level.
@@
 To verify the effectiveness of IEE in reconstructing causal networks, we calculated the Area Under Curve (AUC) values based on multiple simulated time series.
-The AUC value of IEE was $0.871\pm 0.008$, significantly higher than conventional ObsDC indices, such as GC with $0.707\pm 0.069$, TE with $0.816\pm 0.018$, and CCM with $0.837\pm 0.012$.
-The boxplots for AUC values are displayed in Fig.
-3(B).
-AUCs for IEE exhibited a higher mean and a smaller variation compared to the others.
-Additionally, fig.
-S4 presents receiver operating characteristic (ROC) curves.
+The AUC value (mean $\pm$ standard deviation) of IEE was 0.871±0.0080.871$\pm$ 0.0080.871 ± 0.008, significantly higher than conventional ConDC indices, such as GC with 0.707±0.0690.707$\pm$ 0.0690.707 ± 0.069, TE with 0.816±0.0180.816$\pm$ 0.0180.816 ± 0.018, and CCM with 0.837±0.0120.837$\pm$ 0.0120.837 ± 0.012 (boxplots in Fig.
+3(B)).
+The AUCs of IEE exhibited both a higher mean and a lower variance relative to other approaches.
+Additionally, receiver operating characteristic (ROC) curves for the four indices were presented in fig.
+S4.
 To show the capability for ranking the importance of causal influences, we conducted two tests.
@@
 IEE accurately discerned the IntDCs, with Node 6 showing the strongest value.
-A comparison between IEE for the IntDC and the results from GC, TE, and CCM for the ObsDC can be found in figs.
-S5 and S6.
-GC and TE failed to rank the influence from Node 1 accurately (fig.
-S5), while CCM suffered from the false-positive causal detection from Node 8 to Node 7 (fig.
+A comparison showed that GC and TE failed to rank the influence from Node 1 accurately in the first test (fig.
+S5), while CCM suffered from the false-positive causal detection from Node 8 to Node 7 in the second test (fig.
 S6).
@@
 (B) presents a boxplot of the Area Under Curve (AUC) values for the network reconstruction.
-IEE outperforms GC, TE, and CCM significantly in accuracy and stability.
+IEE (0.871±0.0080.871$\pm$ 0.0080.871 ± 0.008) outperforms GC (0.707±0.0690.707$\pm$ 0.0690.707 ± 0.069), TE (0.816±0.0180.816$\pm$ 0.0180.816 ± 0.018), and CCM (0.837±0.0120.837$\pm$ 0.0120.837 ± 0.012) significantly in accuracy and stability.
 (C) demonstrates the IEE from Node 1 to the other nine nodes, while (D) shows the IEE received by Node 7 from the other nodes.
-IEE accurately captures the interventional flow and IntDC in cascade topology.
-IEE can quantify IntDC effectively without requiring interventions or perturbations
+IEE accurately captures the interventional information flow and IntDC in the cascade topological network.
+3.4 IEE quantifies IntDC without requiring additional experimental interventions
 IEE is specifically designed for calculating IntDC solely from observed time-series data without requiring additional perturbations to the system.
 To validate its accuracy, we conducted a comparison against true perturbed deviations using chaotic neural networks (chNNs) as a model system (?, ?).
-The chNN comprises an output variable $\bm{x}=(x_{1},x_{2},\dots,x_{N_{\text{node}}})^{T}$ and two internal variables, with $N_{\text{node}}$ representing the number of neurons in the network.
+The chNN comprises an output variable $\linebreak\bm{x}=(x_{1},x_{2},\dots,x_{N_{\text{node}}})^{T}$ and two internal variables, with $N_{\text{node}}$ representing the number of neurons in the network.
 We chose $N_{\text{node}}=10$ in simulations.
@@
 to quantify the true influence of the intervention, i.e.
-IntDC, from $i$ to $j$.
-By conducting $100$ randomly simulated chNNs with interventions (10 removed nodes for each chNN), we recorded $1000$ KLD values and compared them with the results of IEE (calculated from $\bm{x}_{\text{obs}}$ solely).
-In Fig.
-4(A), gray dots are the $1000$ samples of causal edges between neurons, while the red line stands for the linear regression.
-We observed that $\text{IEE}[x_{i}\rightarrow x_{j}]$ exhibited a positive linear correlation with the KLD $D_{ij}$ (with $R^{2}=0.869$).
-Results for GC (with $R^{2}=0.572$), TE (with $R^{2}=0.912$), and CCM (with $R^{2}=0.622$) are shown in Figs.
-4(B-D), which imply that GC and CCM cannot linearly reflect KLD.
-Furthermore, the violin plots of cosine similarity $S_{c}$ (see SM) between KLD and the four causal indices on $100$ chNNs are displayed in Fig.
-4(E).
-IEE ($0.904\pm 0.024$) had the highest mean similarity with KLD, compared to GC ($0.711\pm 0.076$), TE ($0.888\pm 0.022$), and CCM ($0.687\pm 0.033$).
-These results indicate that IEE can accurately quantify IntDC from $\bm{x}_{\text{obs}}$ alone, alleviating the need for additional interventions to the dynamics.
+ground truth of the IntDC, from $i$ to $j$.
+Liang’s Information flow adopts a similar intervention concept by treating a variable as fixed  (?, ?).
+By conducting 100100100 randomly simulated chNNs with interventions (in each simulation, one node is removed from each chNN with $N_{\text{node}}$ nodes), we recorded 100010001000 KLD values and compared them with the results of IEE (calculated from $\bm{x}_{\text{obs}}$ solely).
+We observed that $\text{IEE}[x_{i}\rightarrow x_{j}]$ exhibited a positive linear correlation with the KLD $D_{ij}$ (with $R^{2}=0.869$) in Fig.
+4(A), where gray dots are the 100010001000 samples of causal edges between neurons and the red line stands for the linear regression.
+Results for GC (with $R^{2}=0.572$), TE (with $R^{2}=0.912$), and CCM (with $R^{2}=0.622$) indicated that GC and CCM could not linearly reflect KLD (Figs.
+4(B-D)).
+Furthermore, we drew the violin plots of cosine similarity $S_{c}$ (see SM) between KLD and the four causal indices on 100100100 chNNs (Fig.
+4(E)).
+IEE (0.904±0.0240.904$\pm$ 0.0240.904 ± 0.024) had the highest mean similarity with KLD, compared to GC (0.711±0.0760.711$\pm$ 0.0760.711 ± 0.076), TE (0.888±0.0220.888$\pm$ 0.0220.888 ± 0.022), and CCM (0.687±0.0330.687$\pm$ 0.0330.687 ± 0.033).
+These results present that IEE can accurately quantify IntDC from $\bm{x}_{\text{obs}}$ alone, alleviating the need for additional interventions to the dynamics.
 Figure 4: The performance of IEE when measuring the influence of perturbations on the chaotic neural networks (chNNs).
 The Kullback-Leibler divergence (KLD) is utilized to measure the true influence or IntDC between neurons.
-In (A-D) , $1000$ samples for different perturbations on chNNs are represented by gray dots, along with the linear regressions (the red lines) between KLD and IEE/GC/TE/CCM, respectively.
+In (A-D) , 100010001000 samples for different perturbations on chNNs are represented by gray dots, along with the linear regressions (the red lines) between KLD and IEE/GC/TE/CCM, respectively.
 (E) displays the violin plots of cosine similarity between KLD and IEE/GC/TE/CCM.
-The mean values and standard variations of the cosine similarity are IEE ($0.904\pm 0.024$), GC ($0.711\pm 0.076$), TE ($0.888\pm 0.022$), and CCM ($0.687\pm 0.033$).
+The mean values and standard deviations of the cosine similarity are IEE (0.904±0.0240.904$\pm$ 0.0240.904 ± 0.024), GC (0.711±0.0760.711$\pm$ 0.0760.711 ± 0.076), TE (0.888±0.0220.888$\pm$ 0.0220.888 ± 0.022), and CCM (0.687±0.0330.687$\pm$ 0.0330.687 ± 0.033).
 IEE can linearly reflect the KLD with high similarity, indicating its effectiveness in accurately quantifying IntDC without additional perturbed data on chNNs.
-Application of IEE in inferring neural connectomes of C. elegans
+3.5 Application of IEE in inferring neural connectomes of C. elegans
 We applied the IEE criterion to infer the neural connectomes of Caenorhabditis elegans (C.
 elegans), a model organism known for its comprehensively studied nervous system  (?).
-Calcium fluorescence imaging time series data from $31$ neurons with specific functions of a freely moving C.
+Calcium fluorescence imaging time series data from 313131 neurons with specific functions of a freely moving C.
 elegans were collected from Kato et al.
@@
 S7).
-The Fig.
-5(C) shows the cosine similarity matrix between different neurons.
-We clustered the neurons into $7$ clusters after data preprocessing (see details in Supplementary Text and table S5).
-Figure 5(D) presents the ground truth of neural connectomes in a directed graph, which was detected by electron microscopy in references  (?, ?, ?).
-We calculated the IEE between different clusters to infer the IntDC network.
-The ROC curve of IEE is shown in Fig.
-5(F) with an AUC value $0.882$.
-The maximum Youden index, maximum concordance probability, and minimum distance to the point $(0,1)$ gave the same optimal operating point (OOP).
-At the OOP, Fig.
-5(E) displays the inferred connectomic network by IEE.
-Comparisons with other ObsDC indices, i.e.
-GC/TE/CCM, revealed that IEE exhibited the highest AUC value, best OOP, largest similarity to the ground truth, lowest false positive at the OOP, and lowest false negative at the OOP (see Table 2, figs.
+The cosine similarity between different neurons were calculated and the neurons were clustered into 777 clusters after data preprocessing (see Fig.
+5(C), Supplementary Text and table S5).
+We presented the ground truth of neural connectomes between clusters in a directed graph (Fig.
+5(D)), which was detected by electron microscopy  (?, ?, ?).
+The IEE values between different clusters were calculated to infer the IntDC network.
+The ROC curve of IEE got a high AUC value 0.8820.8820.882 (Fig.
+5(F)).
+By maximum Youden index, maximum concordance probability, and minimum distance to the point (0,1)(0,1)( 0 , 1 ), IEE gave the same optimal operating point (OOP).
+At the OOP, IEE provided an inferred connectomic network with only 1 false positive and 4 false negatives (Fig.
+5(E)).
+We also compared IEE with other ConDC indices, i.e.
+GC/TE/CCM, which revealed that IEE exhibited the highest AUC value, best OOP, largest similarity to the ground truth, lowest false positive at the OOP, and lowest false negative at the OOP (see Table 2, figs.
 S8-S10, and table S6).
-These findings demonstrate the efficacy of IEE in reconstructing the neural connectomes of C.
-elegans.
-Figure 5: Application of the IEE on inferring the neural connectomes of C.
-elegans.
-(A) is an illustration of C.
-elegans.
-(B) displays the calcium imaging time series of $31$ key neurons, with their names listed on the left and colors representing different clusters.
-(C) is the cosine similarity matrix between neurons, aiding the clustering process.
-(D) is the true connectomes between $7$ clusters, while (E) is the inferred causal network at the the optimal operating point (OOP) determined by IEE.
-The red edge represents a false positive, and the black dashed edges represent false negatives.
-(F) shows the ROC curve (the blue line) for IEE.
-The same OOP is obtained under three criteria, i.e.
-the maximum Youden index (the blue circle), the maximum concordance probability (the green dot), and the minimum distance to the point $(0,1)$ (the red cross).
-The shaded area around the ROC curve represents the $95\%$ confidence interval obtained through bootstrapping.
-The AUC value of IEE is $0.882$.
-Application of IEE to COVID-19 transmission in Japan
+Additionally, we inferred the connectomic network using the PCMCI algorithm with two approaches: the partial correlation (PCMCI-ParCorr) and CMI test (PCMCI-CMI).
+PCMCI used an independence test with the only parameter being the significance level $\alpha_{\text{PC}}$.
+The false positive edges for PCMCI-ParCorr ($\alpha_{\text{PC}}=0.05$) and PCMCI-CMI ($\alpha_{\text{PC}}=0.01$) are 555 and 666, respectively.
+Both approaches have 444 false negative edges.
+A detailed description of the procedure and results can be found in the Supplementary Text, figs.
+S10(e-f) and table S6.
+These findings demonstrate the efficacy of IEE in reconstructing the neural connectomes of C. elegans.
+[…]
+3.6 Application of IEE to COVID-19 transmission in Japan
 We validated that IEE was a promising indicator for assessing the transmission dynamics of infectious diseases.
-Our study collected the daily confirmed COVID-19 cases from all $47$ prefectures in Japan, spanning a duration of 1209 days from January 16, 2020 to May 8, 2023 (fig.
+Our study collected the daily confirmed COVID-19 cases from all 474747 prefectures in Japan, spanning a duration of 1209 days from January 16, 2020 to May 8, 2023 (fig.
 S11).
 First, we calculated the IEE indices for IntDC among different prefectures.
-IEE effectively ranked the influence across prefectures, with the top five affected areas by Tokyo being Kanagawa, Chiba, Saitama, Aichi, and Osaka.
+IEE effectively ranked the influence across prefectures, with the top five affected areas by Tokyo being Kanagawa (with IEE value 0.6010.6010.601), Chiba (with IEE value 0.5680.5680.568), Saitama (with IEE value 0.5340.5340.534), Aichi (with IEE value 0.4530.4530.453), and Osaka (with IEE value 0.4350.4350.435).
 Remarkably, this ranking coincides with the geographical proximity and socio-economic connections to Tokyo.
@@
 The element $D_{ij}^{\text{COVID}}$ served as a benchmark for quantifying COVID-19 transmission from prefecture $i$ to $j$.
-The IEE values demonstrated a strong linear correlation with $\ln D^{\text{COVID}}$, particularly when assessing the causality from Tokyo/Osaka to other prefectures (Figs.
+The IEE values demonstrated a strong linear correlation with $\ln D^{\text{COVID}}$, particularly when assessing the causality from Tokyo/Osaka to other prefectures (with Pearson correlation coefficients, i.e.
+PCCs, 0.906-0.906- 0.906 for Tokyo and 0.860-0.860- 0.860 for Osaka, Figs.
 6(A-B)).
-This correlation remained significant for other regions such as Aichi, Hokkaido, Fukuoka, and Okinawa (see fig.
-S12).
-Further, IEE outperformed ObsDC indices in practicality.
-Notably, the absolute value of the Pearson correlation coefficient (PCC) between IEE and $\ln D^{\text{COVID}}$ was $0.724$ substantially higher than those obtained by GC $0.161$, TE $0.284$, and CCM $0.568$ (average over specific prefecture, see the second to last column in Table 3 and the boxplot in Fig.
+This correlation remained high for other regions (see fig.
+S12), such as Aichi (with PCC value 0.811-0.811- 0.811), Hokkaido (with PCC value 0.751-0.751- 0.751), Fukuoka (with PCC value 0.752-0.752- 0.752), and Okinawa (with PCC value 0.811-0.811- 0.811).
+Further, IEE outperformed ConDC indices in practicality.
+Notably, the absolute value of the PCC between IEE and $\ln D^{\text{COVID}}$ was 0.7240.7240.724 substantially higher than those obtained by GC 0.1610.1610.161, TE 0.2840.2840.284, and CCM 0.5680.5680.568 (average over specific prefecture, see the second to last column in Table 3 and the boxplot in Fig.
 6(C)).
-Table 3 presented more detailed comparisons to support the effectiveness of IEE (with Tokyo for specific illustration in fig.
+More detailed comparisons to support the effectiveness of IEE were presented in Table 3 (with specific illustration for Tokyo in fig.
 S13).
@@
 (C) is the boxplot for $|\text{PCC}(c_{ij}^{\text{Inf}},\ln D_{ij}^{\text{COVID}})|$ with average over fixed $i$, where PCC is the Pearson correlation coefficient and $c^{\text{Inf}}$ is the inferred causal strengths by IEE, GC, TE, or CCM.
-The PCC linearity of IEE ($0.724\pm 0.116$) is significantly higher than that of GC ($0.161\pm 0.099$), TE($0.284\pm 0.168$), and CCM ($0.568\pm 0.174$).
+The PCC linearity of IEE (0.724±0.1160.724$\pm$ 0.1160.724 ± 0.116) is significantly higher than that of GC (0.161±0.0990.161$\pm$ 0.0990.161 ± 0.099), TE(0.284±0.1680.284$\pm$ 0.1680.284 ± 0.168), and CCM (0.568±0.1740.568$\pm$ 0.1740.568 ± 0.174).
 (D) shows the GRN surrounding Clock, a key circadian gene.
-(E) demonstrates the ROC curves and AUC values of the IntDC index IEE, and three ObsDC indices GC/TE/CCM.
-Application of IEE to investigating circadian rhythms
+(E) demonstrates the ROC curves and AUC values of the IntDC index IEE, and three ConDC indices GC/TE/CCM.
+3.7 Application of IEE to investigating circadian rhythms
 We investigated the gene regulatory networks (GRNs) related to key genes on the circadian rhythm.
@@
 Figure 6(D) displayed the GRN around Clock at the protein level, where the transcription factor Clock is phosphorylated by PFK family genes.
-The ROC curves and AUC values for IEE/GC/TE/CCM were showed in Fig.
-6(E).
-IEE designed for quantifying IntDC presented a higher AUC value ($0.737$) than ObsDC indices ($0.490$ for GC, $0.610$ for TE, and $0.644$ for CCM).
-In fig.
-S14, we also tested the indices on the GRN surrounding Cry1/Cry2, where IEE outperformed the others.
-Concluding remarks and perspective
+We calculated the IEE/GC/TE/CCM to reconstruct the GRN (ROC curves and AUC values in Fig.
+6(E)).
+IEE designed for quantifying IntDC presented a higher AUC value (0.7370.7370.737) than ConDC indices (0.4900.4900.490 for GC, 0.6100.6100.610 for TE, and 0.6440.6440.644 for CCM).
+We also tested the indices on the GRN surrounding Cry1/Cry2 (fig.
+S14), where IEE obtained the highest AUC value (0.6390.6390.639) and outperformed the others (0.4280.4280.428 for GC, 0.6190.6190.619 for TE, and 0.6140.6140.614 for CCM).
+4 Conclusions and discussions
 In summary, we have established a framework of IntDC based on dynamical systems theory and introduced the IEE criterion to quantify IntDC in this study.
-IEE measures the information flow between interventional causes and effects in the delay-embedding space, making it suitable for analyzing non-linear and non-separable systems (?, ?).
+IEE measures the information flow between interventional causes and effects in the delay-embedding space, making it suitable for analyzing non-linear and non-separable systems (?, ?, ?).
 Moreover, IEE is able to infer IntDC solely from observational time-series data, without requiring additional perturbations to the system.
 This property makes IEE particularly suitable for non-intervention systems.
-Both theoretical derivations and numerical examples presented in this study provide strong evidence to support the high effectiveness, accuracy and robustness of IEE in detecting IntDC and reconstructing networks.
+Both theoretical derivations and numerical examples presented in this study provide strong evidence to support the effectiveness, accuracy and robustness of IEE in detecting IntDC and reconstructing networks.
 Furthermore, IEE serves as an effective tool for evaluating and ranking the causal dependence between variables within a dynamical system.
 Through real-world examples, we illustrated the promising applicability of IEE in diverse fields such as regulatory inference and disease transmission studies.
-There are several compelling issues that warrant further investigation and future development.
+Information-flow-based methods have been developing rapidly in recent years  (?, ?, ?).
+Liang’s information flow theory, albeit with full nonlinearity in its theoretical formalism, is yet to be implemented for practical applications on real-world data.
+Existing implementations still rely on assumptions such as independent white noise and linearity when evaluating information flow from data.
+Our IEE algorithm, inspired by the conception information flow, is specifically designed to handle fully nonlinear systems and has demonstrated applicability across a variety of real datasets.
+The proposed IntDC framework, together with the IEE algorithm, advances the frontier of causality research by providing a practical and effective tool for nonlinear causal inference.
+Runge et al.
+(?) summarized recent advances and key challenges in the field of causal inference and developed the platform https://causeme.net, which serves as a valuable resource for researchers.
+Causal inference holds great promise for applications in earth system science and beyond.
+Our proposed IEE algorithm, together with the concept of IntDC, provides a novel and practical approach for investigating causality in complex nonlinear systems.
+There are several important open issues that warrant further investigation and future development.
 (1) The theory of IntDC and the IEE criterion can be extended to analyze causal relations between groups of variables.
 Specifically, in Eq.
-(6), the variables $x$ and $y$ can be generalized to high-dimensional vectors, enabling the investigation of complex interactions among multiple variables.
-(2) The data in computing IEE can be non-uniformly sampled in the temporal space.
-We used the time-delayed embedding Eq.
+(2), the variables $x$ and $y$ can be generalized to multivariate vectors, enabling the investigation of complex interactions among multiple variables.
+(2) The data used for computing IEE can be non-uniformly sampled in time.
+In this study, we employed the standard time-delayed embedding Eq.
 (3) and Eq.
-(3) in this paper, while according to the general Takens’ embedding theorem, any $L+1$ observations of $x_{t}$ and $y_{t}$ can form a topological diffeomorphism to $\mathcal{M}_{X}$ and $\mathcal{M}_{Y}$, respectively.
-Equation (6) shows that only neighbors in the embedding space matter.
-Thus, IEE algorithm can be extended to infer causality using diverse datasets sampled at different intervals.
-(3) To distinguish direct and indirect causality, we can generalize IEE to its conditional version
-$\text{cIEE}[x\rightarrow y|z]=\text{CMI}(\delta\bm{X}_{t},\delta\bm{Y}_{t+1}|\bm{Y}_{t+1},\bm{Z}_{t},\delta\bm{Z}_{t}),$ (10)
-where “cIEE” is short for the conditional IEE, $z$ is a third variable whose time-delay vector is $\bm{Z}_{t}$, and $\delta\bm{Z}_{t}$ represents the intervened deviation on $\bm{Z}_{t}$.
+(4).
+However, according to the generalized Takens’ embedding theorem, any set of $L+1$ observations of $x_{t}$ and $y_{t}$ can construct topological diffeomorphisms to the manifolds $\mathcal{M}_{X}$ and $\mathcal{M}_{Y}$, respectively.
+As shown in Eq.
+(7), only the local neighborhood structure in the embedding space is essential for the computation.
+Therefore, the IEE algorithm can be naturally extended to infer causality from datasets sampled at varying or irregular time intervals.
+(3) To further distinguish direct and indirect causality, we can generalize IEE to its conditional version
+$\text{cIEE}[x\rightarrow y|z]:=\text{CMI}(\delta\bm{X}_{t},\delta\bm{Y}_{t+1}|\bm{Y}_{t+1},\bm{Z}_{t},\delta\bm{Z}_{t}),$ (11)
+where “cIEE” is short for the conditional IEE, $z$ is a third variable whose time-delayed vector is $\bm{Z}_{t}$, and $\delta\bm{Z}_{t}$ represents the intervened deviation on $\bm{Z}_{t}$.
 Equation (11) measures the direct IntDC from $x$ to $y$ conditioned on $z$.
-Further, together with the well-known PC algorithm  (?), cIEE is one feasible way to remove high-order indirect causal edges iteratively and get the direct causal network.
-(4) Moreover, causal inference with deep learning technology represents a frontier in research.
-Notable works include CausalEGM  (?) and Causalformer (?).
-Combining IntDC and neural networks is a promising field.
-Causality theory brings interpretability for artificial intelligence, while big model can provide new discoveries in causal relationships.
-We are now entering a brilliant “big causality era”.
+Further, in combination with the well-known PC algorithm  (?) or PCMCI (?), cIEE offers a feasible way to remove high-order indirect causal edges iteratively and reconstruct the direct causal network.
+(4) In  (?, ?, ?), researchers have pointed out that time-delayed embedding algorithms are particularly suitable for deterministic dynamical systems with an attractive manifold.
+However, the coupling strength between variables could influence the accuracy of these methods.
+Stark et al.
+(?) extended Takens’ embedding to stochastic systems, providing a theoretical basis for applying embedding approaches beyond purely deterministic dynamics.
+In the Results section of this study, we demonstrated the effectiveness of IEE for Logistic systems with Gaussian noises.
+Nevertheless, how systematic factors, such as coupling strength and noise characteristics, affect the quantification of causal strength in more general settings remains an open question.
+(5) Moreover, causal inference empowered by deep learning is emerging as a frontier in research.
+Notable advances include CausalEGM  (?), Causalformer (?), intervened reservoir computing (?), and reservoir cross mapping (?).
+Integrating the IntDC framework with neural networks presents a promising direction for further exploration.
+On one hand, causality theory provides interpretability to artificial intelligence; on the other hand, large AI foundation models offer unprecedented capacity to uncover causal relationships in complex data .
+These developments signal the advent of a transformative “big causality era”, in which data-driven discovery and causal reasoning evolve in synergy.
 Acknowledgments
 Funding:
-This work was supported by the National Key R&D Program of China 2022YFC2704604, the National Natural Science Foundation of China No.
-12301620 and No.
-32100509, the Moonshot R&D Grant Number JPMJMS2021, Institute of AI and Beyond of UTokyo, the International Research Center for Neurointelligence (WPI-IRCN) at The University of Tokyo Institutes for Advanced Study (UTIAS), JSPS KAKENHI Grant Number JP20H05921, the Cross-ministerial Strategic Innovation Promotion Program (SIP), and the 3rd period of SIP “Smart energy management system” Grant Number JPJ012207.
-Author contributions:
-Conceptualization: J.S., K.A., L.C., and W.L.
-Methodology: J.S., K.A., L.C., and W.L.
-Investigation: J.S., Y.L., J.Z., and S.L.
-Visualization: J.S., Y.L., J.Z., S.L.
-Supervision: K.A., L.C., and W.L.
-Writing – original draft: J.S, Y.L., J.Z., and L.C.
-Writing – review & editing: All authors.
+This work was supported by the National Natural Science Foundation of China [grant numbers 42450192, 42450135, 12301620, 11925103 and 32100509], the National Key R&D Program of China [grant number 2022YFC2704604], the Shanghai Municipal Data Bureau special fund for urban digital transformation [grant number 202401065], the AI for Science Foundation of Fudan University [grant number FudanX24AI041], the AI for Science Program of Shanghai Municipal Education Commission [grant number 24RGZN10], the Japan Science and Technology Agency Moonshot R&D [grant number JPMJMS2021], the Institute of AI and Beyond of UTokyo, the International Research Center for Neurointelligence (WPI-IRCN) at The University of Tokyo Institutes for Advanced Study (UTIAS), the Cross-ministerial Strategic Innovation Promotion Program (SIP), the 3rd period of SIP [grant number JPJ012207].
 Competing interests: